### v3

# Data-Locality-Aware Task Assignment and Scheduling for Distributed Job Executions

Abstract
This paper investigates a data-locality-aware task assignment and scheduling problem aimed at minimizing job completion times for distributed job executions. Without prior knowledge of future job arrivals, we propose an optimal balanced task assignment algorithm (OBTA) that minimizes the completion time of each arriving job. We significantly reduce OBTA’s computational overhead by narrowing the search space of potential solutions. Additionally, we extend an approximate algorithm known as water-filling (WF) and nontrivially prove that its approximation factor equals the number of task groups in the job assignment. We also design a novel heuristic, replica-deletion (RD), which outperforms WF. To further reduce the completion time of each job, we expand the problem to include job reordering, where we adjust the order of outstanding jobs following the shortest-estimated-time-first policy. Extensive trace-driven evaluations validate the performance and efficiency of the proposed algorithms.
Index Terms:
task assignment, job scheduling, distributed job execution, approximate analysis, trace-driven evaluation.
I Introduction
Scheduling plays a critical role in the performance of big data analysis and high-performance computing, where jobs are executed across multiple distributed servers [1, 2]. These jobs typically consist of numerous tasks, each requiring access to distinct data chunks that might be replicated across different locations. Data locality, which involves assigning tasks to servers holding the required data chunks, is crucial in this context [3]. Designing online scheduling algorithms that uphold data locality in distributed job executions, without prior knowledge of job arrivals, involves two key aspects: task assignment and job reordering. Task assignment involves allocating tasks within a job to appropriate servers, conceptualized as a semi-matching problem in a bipartite graph where tasks and servers are distinct sets of nodes. Job reordering addresses the sequence in which the arrived but unfinished job tasks are executed, facilitating optimized task assignments and reducing overall job completion times.
Prior studies have addressed these issues either in isolation or in tandem [4, 5, 6, 7]. For instance, Hung et al. [4] proposed a Workload-Aware Greedy Scheduling (SWAG) strategy that schedules jobs based on their estimated completion times, prioritizing the job with the minimal completion time. Guan et al. [5] presented an offline fairness-aware task assignment algorithm that ensures max-min fairness in resource allocation across all servers. However, both [4] and [5] did not consider data replication across multiple servers, significantly reducing the complexity of the scheduling problem. Beaumont et al. [6] investigated data replication-aware scheduling, but their study focused on a single job and did not consider job reordering. Additionally, their proposed semi-matching-based strategies are only applicable to homogeneous servers with the same processing capacity and homogeneous tasks with the same execution time, which oversimplifies real-world scenarios. Guan et al. [7] studied an online scheduling problem for distributed job executions where the data chunks requested by each job could be available at multiple servers. They solved the task assignment problem using a well-constructed flow network, mapping the assignment of tasks into a maximum flow problem. For job reordering, they adopted a similar approach to [4], considering job priorities. However, they did not provide an in-depth performance analysis of the proposed algorithms.
Our work primarily focuses on the task assignment aspect of scheduling while maintaining data locality. This challenge is addressed in an online setting, where the arrival of jobs over time demands immediate scheduling decisions. The objective is to minimize the completion time of all arrived jobs, defined as the time duration from a job’s arrival to the completion of all its tasks. We study the problem in two scenarios. In the first scenario, the outstanding jobs are pending in the queues in a first-in-first-out (FIFO) manner. In the second scenario, the outstanding jobs can be prioritized and then reordered. Our contributions are summarized as follows.
  • We formulate the task assignment problem as a bipartite graph-based non-linear program and develop an optimal algorithm called OBTA to solve it efficiently by significantly narrowing the search space of potential solutions.
  • We extend an approximate algorithm called water-filling (WF), initially proposed for homogeneous scenarios in [7], to our settings. We nontrivially prove that WF is $K$-approximate, where $K$ is the number of task groups in a job.
  • A novel heuristic, replica-deletion (RD), is developed for task assignment. RD assigns the tasks of a job to servers with a complexity of $\mathcal{O}(M^{2}\cdot n\log n)$, where $M$ is the number of servers and $n$ is the number of tasks in a job.
  • We extend WF to accommodate job reordering and introduce an early-exit technique to expedite the process, resulting in the OCWF-ACC (accelerated order-conscious WF) algorithm.
  • Our trace-driven simulations underscore the effectiveness and efficiency of the proposed algorithms. OBTA, in particular, demonstrates a notable reduction in computation overhead, while WF offers a low-overhead yet effective approximation. RD further enhances performance. Additionally, compared to a job-reordered task assignment algorithm without early-exit, OCWF-ACC significantly reduces computation overhead.
The rest of this paper is organized as follows. Sec. II presents the problem formulation. Sec. III introduces the task assignment algorithms for FIFO queues. Sec. IV expands the algorithms to include job reordering. Experimental results using real job traces are discussed in Sec. V. Sec. VI reviews the related work, and Sec. VII concludes the paper.
II Problem Formulation
We consider a distributed computing system consisting of $M$ servers, indexed by $m$ for $m\in\mathcal{M}:=\{1,...,M\}$. The system stores a collection of items, such as key-value pairs, objects, files, etc. They are partitioned into equally-sized data chunks. Each server holds some data chunks, and a data chunk can be replicated across multiple servers. For each $m\in\mathcal{M}$, we denote by $\mathcal{D}_{m}$ the set of data chunks stored in server $m$. The intersection $\mathcal{D}_{m}\cap\mathcal{D}_{m^{\prime}}$ can be non-empty for different $m$ and $m^{\prime}$ in $\mathcal{M}$. We make no particular assumptions about data chunk distributions. We assume that the placement of data chunks is given and does not change over time.
Jobs arrive in the system online, one by one, for execution. Each job consists of multiple independent tasks that process different data partitions. Each task takes one data chunk as input, and the tasks of the same job are assumed to have equal demands on the compute resources. We denote by $c\in\mathbb{N}^{+}$ the $c$-th job (also referred to as job $c$) in chronological order. Further, we use $\mathcal{T}_{c}$ to represent the set of tasks in job $c$, and $d_{r}$ to represent the data chunk demanded by a task $r\in\mathcal{T}_{c}$. For each task $r\in\mathcal{T}_{c}$, we define
$\mathcal{S}^{r}:=\big{\{}m\in\mathcal{M}\mid d_{r}\in\mathcal{D}_{m}\big{\}}$ (1)
as the set of its available servers. Due to data locality, each task should be assigned to one of its available servers for processing. We assume that each server $m$ has a sufficiently large queue, denoted by $q_{m}$, to store the outstanding job tasks. For analysis purposes, we divide the time into time slots of identical length.
We denote by $\mu_{m}^{c}$ the profiled number of tasks for job $c$ that can be processed at server $m$ within a time slot. We define the initial busy time of each server $m$ just before job $c$’s arrival, denoted by $b_{m}^{c}$, as the number of time slots required to finish all the outstanding job tasks in the corresponding queue. The initial busy time $b_{m}^{c}$ can be estimated using the following equation:
$b_{m}^{c}=\sum_{h=1}^{c-1}\left\lceil\frac{o_{m}^{h}}{\mu_{m}^{h}}\right\rceil,$ (2)
where $o_{m}^{h}\leq|\mathcal{T}_{h}|$ is the number of outstanding tasks from job $h$ queued in $q_{m}$ just before job $c$’s arrival.
Our task is to make assignment and scheduling decisions for each arriving job. The objective is to minimize the completion time of each job. For each job, the completion time is defined as the time duration between its arrival and the completion of its last processed task. This objective is equivalent to minimizing the completion time of the last processed task issued by the job. We study the problem in two scenarios:
  • FIFO queues. The outstanding job tasks are queued in servers in a FIFO manner. The key is to assign the tasks of each arriving job to their available servers in a way that minimizes the job completion time by keeping the estimated busy times of the servers as balanced as possible.
  • Prioritized reordering. The outstanding job tasks in each queue can be reordered. The challenge lies in prioritizing the outstanding jobs in such a way that the completion times of the jobs can be further reduced.
We make no particular assumptions about job arrival times.
III Algorithms for FIFO Queues
III-A Matching-based Non-Linear Programming
III-A1 Programming in a Bipartite Graph
When a new job $c$ arrives, we need to assign all the tasks in $\mathcal{T}_{c}$ to appropriate available servers.
We do this by dividing these tasks into task groups, where each task group contains tasks that share the same set of available servers. We denote by $K_{c}$ the number of task groups in job $c$. Further, for each $k\in\mathcal{K}_{c}:=\{1,\ldots,K_{c}\}$, we denote by $\mathcal{T}_{c}^{k}$ and $\mathcal{S}_{c}^{k}$ the set of tasks in the $k$-th group (also referred to as group $k$) and the set of available servers for tasks in the $k$-th group, respectively. By definition, we have
$\mathcal{T}_{c}^{k}=\left\{r\in\mathcal{T}_{c}\mid\mathcal{S}^{r}=\mathcal{S}_{c}^{k}\right\}.$ (3)
The task assignment problem can be modeled as a matching problem with respect to a bipartite graph. The left-side nodes in the bipartite graph are the $K_{c}$ task groups, while the right-side nodes are the $M$ servers. For each $k\in\mathcal{K}_{c}$, an edge $(k,m)$ exists if $m\in\mathcal{S}_{c}^{k}$. We can formulate the assignment problem as a non-linear integer program as follows:
$\mathcal{P}:\qquad\qquad\qquad\min_{\Phi_{c},n_{m}^{k}\in\mathbb{N}^{+}}\Phi_{c}$
s.t. $\left\{\begin{array}[]{ll}\sum_{k\in\mathcal{K}_{c}}n_{m}^{k}\leq\max\left\{\Phi_{c}-b_{m}^{c},0\right\}&\forall m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k},\\ \sum_{m\in\mathcal{S}_{c}^{k}}n_{m}^{k}\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|&\forall k\in\mathcal{K}_{c}.\end{array}\right.$ (6)
Recall that $b_{m}^{c}$ is the initial estimated busy time of server $m$ before the assignment of tasks in $\mathcal{T}_{c}$, calculated by (2). In $\mathcal{P}$, $n_{m}^{k}$ is the number of time slots required to process the tasks in group $\mathcal{T}_{c}^{k}$ at server $m$, given the profiled processing capacity $\mu_{m}^{c}$. Each $n_{m}^{k}$ is associated with an edge $(k,m)$ in the bipartite graph. $\Phi_{c}$ is the estimated completion time of job $c$. The first constraint in (6) means that any available server whose initial estimated busy time is larger than $\Phi_{c}$ should not participate in the assignment of tasks in $\mathcal{T}_{c}$. This constraint ensures balance among the participating servers. The second constraint in (6) ensures that all the tasks in $\mathcal{T}_{c}$ can be processed. The program $\mathcal{P}$ can be solved optimally by commercial solvers, e.g., CPLEX. [^1] https://www.ibm.com/products/ilog-cplex-optimization-studio
III-A2 Narrowing the Search of $\Phi_{c}$
A commercial solver could be slow to solve the non-linear integer program in large-scale scenarios. To reduce the computational overhead, we narrow the search space of $\Phi_{c}$ by analyzing the characteristics of the problem.
After job $c$’s arrival, the number of time slots required to process all the backlogged tasks and the newly assigned tasks issued by job $c$ is upper bounded by
$\Phi_{c}^{+}:=\max_{m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}}\left\{\left\lceil\frac{\sum_{k\in\mathcal{K}_{c}:m\in\mathcal{S}_{c}^{k}}|\mathcal{T}_{c}^{k}|}{\mu_{m}^{c}}\right\rceil+b_{m}^{c}\right\}.$ (7)
$\Phi_{c}^{+}$ is obtained by assuming that for each available server $m$, all the tasks in $\mathcal{T}_{c}$ are assigned to it. Correspondingly, the lower bound of $\Phi_{c}$ is
$\Phi_{c}^{-}:=\max_{k\in\mathcal{K}_{c}}x_{k},$ (8)
where $x_{k}$ is the minimum integer such that
$\sum_{m\in\mathcal{S}_{c}^{k}}\left(\max\left\{x_{k}-b_{m}^{c},0\right\}\right)\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|.$ (9)
Note that $x_{k}$ is the minimal number of time slots required to process all the tasks in $\mathcal{T}_{c}^{k}$ under the assumption that group $k$ is the only task group of job $c$. With the above analysis, we can reduce the search space of $\Phi_{c}$ from $\mathbb{N}^{+}$ to the integer interval $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$.
III-A3 The Matching-based Algorithm
Figure 1: Divide $\big{[}\Phi_{c}^{-},\Phi_{c}^{+}\big{]}$ into disjoint subranges.
By introducing $\Phi_{c}^{-}$ and $\Phi_{c}^{+}$, the search space of $\Phi_{c}$ is greatly narrowed. However, $\mathcal{P}$ is still non-linear since the first constraint in (6) is piecewise. To solve $\mathcal{P}$ efficiently, we divide $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$ into several disjoint sub-intervals based on the initial estimated busy times of available servers. The division is illustrated in Fig. 1. Specifically, we sort the initial estimated busy time $b_{m}^{c}$ of each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ in ascending order, and we denote the sorted values by $b_{1}^{\prime},b_{2}^{\prime},\ldots$. We then denote by $i$ ($j$) the minimal (maximal) index of these values such that $b_{i}^{\prime}\geq\Phi_{c}^{-}$ ($b_{j}^{\prime}\leq\Phi_{c}^{+}$). Then, we can solve $\mathcal{P}$ with the search space of $\Phi_{c}$ falling into each sub-interval $[\Phi_{c}^{-},b_{i}^{\prime}),[b_{i}^{\prime},b_{i+1}^{\prime}),\ldots,[b_{j}^{\prime},\Phi_{c}^{+}]$ in turn. Each of these problems is transformed into a linear integer program since the first constraint of (6) is no longer piecewise. The sub-intervals are checked one by one. If $\mathcal{P}$ is solvable within the current sub-interval, then its solution is the optimal one since the remaining sub-intervals cannot contain a smaller $\Phi_{c}$.
With $\mathcal{P}$ solved as detailed above, we design a task assignment algorithm named Optimal Balanced Task Assignment (OBTA). The pseudocode is presented in Algorithm 1. When the profiling of $\{\mu_{m}^{c}\}_{m,c}$ is precise, the solution of OBTA is optimal since OBTA outputs the optimal $n_{m}^{k}$’s.
1 while a new job $c$ arrives do
2       for each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
3             Estimate $b_{m}^{c}$ by (2)
4      Solve $\mathcal{P}$ as described in Sec. III-A3, and get the result $\{n_{m}^{k}\}_{m,k}$
5       for each task group $k\in\mathcal{K}_{c}$ do
6             for each server $m\in\mathcal{S}_{c}^{k}$ do
7                   if server $m$ is not the last server in $\mathcal{S}_{c}^{k}$ then
8                         Assign $n_{m}^{k}\cdot\mu_{m}^{c}$ group-$k$ tasks to server $m$
9                   else
10                         Assign all the remaining group-$k$ tasks to server $m$
11
12
13
Algorithm 1 OBTA
III-B Water-Filling
Solving a group of integer programs might still be time-consuming for large-scale distributed computing systems. To improve efficiency, we propose another task assignment algorithm called water-filling (WF), extended from [7], which assigns tasks one group at a time in an approximate manner.
III-B1 Algorithm Design
We denote by $b_{m}^{c}(k)$ the estimated busy time of server $m$ after assigning the tasks in group $k$. By definition, for each $m\in\mathcal{M}$, we have
$b_{m}^{c}(0):=b_{m}^{c}\leq b_{m}^{c}(1)\leq\cdots\leq b_{m}^{c}\left(K_{c}\right).$ (10)
For each $k\in\mathcal{K}_{c}$, WF allocates $\max\{\xi_{k}-b_{m}^{c}(k-1),0\}$ time slots on each available server $m\in\mathcal{S}_{c}^{k}$, where $\xi_{k}$ is the minimal integer such that
$\sum_{m\in\mathcal{S}_{c}^{k}}\left(\max\left\{\xi_{k}-b_{m}^{c}(k-1),0\right\}\right)\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|.$ (11)
The calculation of $\xi_{k}$ is similar to $x_{k}$ in (9). It is the estimated time for all the tasks in group $k$ to be processed by server $m$, given the assignment of tasks in $\mathcal{T}_{c}^{1},\ldots,\mathcal{T}_{c}^{k-1}$. We call $m\in\mathcal{S}_{c}^{k}$ a participating server if $\xi_{k}-b_{m}^{c}(k-1)>0$. We then allocate $(\xi_{k}-b_{m}^{c}(k-1))\cdot\mu_{m}^{c}$ tasks in group $k$ to each participating server $m$ (or all the remaining tasks in group $k$ if server $m$ is the last participating server). After that, we update $b_{m}^{c}(k)$ by
$b_{m}^{c}(k)=\max\left\{b_{m}^{c}(k-1),\xi_{k}\right\}$ (12)
for each $m\in\mathcal{S}_{c}^{k}$. WF repeats this procedure for each task group sequentially, performing like filling water, ensuring that the participating servers have the same estimated busy time. Fig. 2 illustrates the idea of water-filling.
Figure 2: Assign tasks in $\mathcal{T}_{c}^{k}$ to their available servers in a water-filling manner. In (i), only a subset of $\mathcal{S}_{c}^{k}$, including servers $1,...,m^{*}$, participate in the assignment of $\mathcal{T}_{c}^{k}$; In (ii), every server in $\mathcal{S}_{c}^{k}$ participates in the assignment of $\mathcal{T}_{c}^{k}$.
The pseudocode of WF is presented in Algorithm 2. The complexity of WF is $\mathcal{O}\left(\sum_{k\in\mathcal{K}_{c}}|\mathcal{S}_{c}^{k}|\log(|\mathcal{T}_{c}^{k}|)\right)$, since for each group $k$, $\xi_{k}$ can be found through a binary search with an overhead of $\mathcal{O}(\log(|\mathcal{T}_{c}^{k}|))$, and WF needs to walk through all the available servers to verify that the searched $\xi_{k}$ satisfies (11). Note that the complexity of WF can also be written as $\mathcal{O}\left(K_{c}\cdot M\cdot\log|\mathcal{T}_{c}|\right)$.
1 while a new job $c$ arrives do
2       for each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
3             Estimate $b_{m}^{c}$ by (2)
4             $b_{m}^{c}(0)\leftarrow b_{m}^{c}$
5      for each task group $k\in\mathcal{K}_{c}$ do
6             Calculate $\xi_{k}$ as detailed in Sec. III-B1
7             for each server $m\in\mathcal{S}_{c}^{k}$ do
8                   if $b_{m}^{c}(k-1)\geq\xi_{k}$ then
9                         continue
10                  if server $m$ is not the last server in $\mathcal{S}_{c}^{k}$ then
11                         Assign $\left(\xi_{k}-b_{m}^{c}(k-1)\right)\cdot\mu_{m}^{c}$ group-$k$ tasks to server $m$
12                   else
13                         Assign all the remaining group-$k$ tasks to server $m$
14                  Update $b_{m}^{c}(k)$ by (12)
15
16
Algorithm 2 WF
III-B2 Approximation Analysis
In this section, we present the performance analysis of WF. The approximation factor of WF against the optimum is defined as
$\alpha_{\textsc{WF}}:=\max_{I}\frac{\text{WF}(I)}{\text{OPT}(I)},$ (13)
where $I:=I\big{(}c,\{b_{m}^{c}\}_{m}\big{)}$ is an arrival instance of a new job $c$ and the initial estimated busy time of each server just before job $c$’s arrival. $\text{OPT}(I)$ and $\text{WF}(I)$ are respectively the maximum estimated busy times of participating servers after the tasks of job $c$ are assigned by the optimal algorithm and WF.
Figure 3: A visualization of the assignment made by WF for the constructed instance.
Theorem 1.
For each newly arrived job $c$, the approximation factor of WF is at least $K_{c}$.
Proof.
The WF-to-optimal ratio is 1 whenever $\mathcal{T}_{c}$ has only one task group or $\mathcal{S}_{c}^{k}\cap\mathcal{S}_{c}^{k^{\prime}}=\emptyset$ holds for any $k\neq k^{\prime}$. Thus, to get a worst-case ratio, we construct an instance $I$ as illustrated in Fig. 3, where the available servers of different task groups overlap.
For simplicity, we assume $\mu_{m}^{c}=1$ and $b_{m}^{c}=0$ for each server $m$. For each group $k$, the number of its available servers is given by $|\mathcal{S}_{c}^{k}|=\sum_{k^{\prime}=1}^{K_{c}-k+1}\theta^{k^{\prime}}$, where $\theta\geq 2$ is an integer. Moreover,
$\mathcal{S}_{c}^{1}\supset\mathcal{S}_{c}^{2}\supset\cdots\supset\mathcal{S}_{c}^{K_{c}},$ (14)
i.e., the available servers of a higher-indexed group are a subset of those of a lower-indexed group. The number of tasks in group $k$ is given by $|\mathcal{T}_{c}^{k}|=\theta\cdot|\mathcal{S}_{c}^{k}|$. The task assignment made by WF is shown in Fig. 3, where $\theta$ time slots are required to process the tasks in each group.
Figure 4: A visualization of the assignment made by OPT for the constructed instance.
By contrast, in OPT, for each $k<K_{c}$, the tasks in group $k$ are assigned to a subset of its available servers $\mathcal{S}_{c}^{k}\backslash\mathcal{S}_{c}^{k+1}$ only, as illustrated in Fig. 4. Then, the number of time slots needed to process all the tasks in group $k$ is
$\Bigg{\lceil}\frac{\theta\cdot\sum_{k^{\prime}=1}^{K_{c}-k+1}\theta^{k^{\prime}}}{\theta^{K_{c}-k+1}}\Bigg{\rceil}=\theta+2.$ (15)
Additionally, the number of time slots needed to process all the tasks in group $K_{c}$ is $\theta$. Hence, by OPT, the completion time of job $c$ is $\textsc{OPT}(I)=\theta+2$.
Thus, the WF-to-optimal ratio is
$\frac{\textsc{WF}(I)}{\textsc{OPT}(I)}=\frac{K_{c}\cdot\theta}{\theta+2},$ (16)
which can be made arbitrarily close to $K_{c}$ as $\theta$ approaches infinity. ∎
Theorem 1 shows that $K_{c}$ is a lower bound of $\alpha_{\textsc{WF}}$. In the following, we prove that $K_{c}$ is also an upper bound of $\alpha_{\textsc{WF}}$.
Theorem 2.
For each newly arrived job $c$, the approximation factor of WF is at most $K_{c}$.
Proof.
We need to prove that $\textsc{WF}(I)\leq K_{c}\cdot\textsc{OPT}(I)$ holds for any possible arrival instance $I$. Hereafter, we drop $I$ for simplicity and use ALG to represent the algorithm itself or the estimated completion time of job $c$ by applying the algorithm interchangeably (ALG can be OPT, WF, WFk, and OPTk, where the last two notations will be introduced soon).
[…]
$\textsc{WF}_{k}=\max_{m\in\bigcup_{k^{\prime}\leq k}\Omega_{c}^{k^{\prime}}}b_{m}^{c}(k),$ (17)
and for each $k<K_{c}$,
$\textsc{WF}_{k+1}=\max\left\{\textsc{WF}_{k},\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\right\}.$ (18)
Recall that $b_{m}^{c}(k)$ is updated by (12). Fig. 5 visualizes $\textsc{WF}_{k}$.
Figure 5: Visualization of $\textsc{WF}_{k}$. Note that $\textsc{WF}_{k}=\textsc{WF}_{k-1}$ for the constructed example.
Since $\mathcal{T}_{c}^{1}\subseteq\mathcal{T}_{c}$, $\textsc{WF}_{1}\leq\textsc{OPT}$. Note that $\textsc{WF}=\textsc{WF}_{K_{c}}$. If we can prove that for each $k<K_{c}$,
$\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT},$ (19)
then we have
$\textsc{WF}=\textsc{WF}_{1}+\sum_{k<K_{c}}(\textsc{WF}_{k+1}-\textsc{WF}_{k})\leq K_{c}\cdot\textsc{OPT},$ (20)
which is exactly our target. In the following, we will show that (19) indeed holds. To do so, we exploit the optimal assignment of the tasks of a job containing the tasks $\mathcal{T}_{c}^{k+1}$ only. Similar to $\textsc{WF}_{k+1}$, we denote by $\textsc{OPT}_{k+1}$ the number of time slots required to complete the processing of tasks in $\mathcal{T}_{c}^{k+1}$. It is easy to see that
$_{k+1}\leq\textsc{OPT}$ OPT (21)
holds for each $k\in\mathcal{K}_{c}$, since $\mathcal{T}_{c}^{k+1}\subseteq\mathcal{T}_{c}$.
To prove (19), we consider different cases of the assignment of tasks in $\mathcal{T}_{c}^{k+1}$ by WF separately.
Case I. $\mathcal{S}_{c}^{k+1}\cap\bigcup_{k^{\prime}\leq k}\Omega_{c}^{k^{\prime}}=\emptyset$. In this case, the available servers for group $k+1$ do not overlap with the participating servers for groups $1,\ldots,k$. Thus, by the definition of WF, such an assignment must be optimal if a job contains only the tasks $\mathcal{T}_{c}^{k+1}$. Hence, by (21), we have
$\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)=\textsc{OPT}_{k+1}\leq\textsc{OPT}.$ (22)
  • As visualized in Fig. 6(i), if $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$, by (18), we have $\textsc{WF}_{k+1}=\textsc{WF}_{k}$. Thus,
    $\textsc{WF}_{k+1}-\textsc{WF}_{k}=0<\textsc{OPT}.$ (23)
  • As visualized in Fig. 6(ii), if $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$, by (18) and (22), we have
    $\textsc{WF}_{k+1}-\textsc{WF}_{k}$ $<\textsc{WF}_{k+1}$
    $=\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)$
[…]
Figure 6: Visualization of Case I. In (i), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$. In (ii), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$.
Case II. $\mathcal{S}_{c}^{k+1}\cap\bigcup_{k^{\prime}\leq k}\Omega_{c}^{k^{\prime}}\neq\emptyset$, i.e., the available servers of group $k+1$ overlap with the participating servers of lower-indexed groups.
If $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$, we again have $\textsc{WF}_{k+1}=\textsc{WF}_{k}$ by (18) and thus, $\textsc{WF}_{k+1}-\textsc{WF}_{k}=0<\textsc{OPT}$.
If $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$, we have $\textsc{WF}_{k+1}>\textsc{WF}_{k}$. In this case, by definition, each participating server of groups $1,\ldots,k$ must also be a participating server of group $k+1$ by WF (because the estimated busy times of all these servers are smaller than $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)$).
For each $m\in\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1}$ (i.e., server $m$ is available to group $k+1$ but does not participate in $\mathcal{T}_{c}^{k+1}$’s assignment), it can be shown that:
  1. 1.
    $b_{m}^{c}(k)\geq\max_{m^{\prime}\in\Omega_{c}^{k+1}}b_{m^{\prime}}^{c}(k+1)$, and
  2. 2.
    by WF, server $m$ does not participate in the assignment of groups $1,\ldots,k$.
Property 1) follows directly from the definition of WF (otherwise, server $m$ must be a participating server of group $k+1$). Property 2) follows from the above observation that each participating server of groups $1,\ldots,k$ must also be a participating server of group $k+1$ by WF. By properties 1) and 2), for each server $m\in\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1}$, the initial estimated busy time $b_{m}^{c}(0)$ satisfies
$b_{m}^{c}(0)=b_{m}^{c}(k)\geq\max_{m^{\prime}\in\Omega_{c}^{k+1}}b_{m^{\prime}}^{c}(k+1).$ (25)
With this result, we prove $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$ by comparing $\Omega_{c}^{k+1}$ and $\beth_{c}^{k+1}$, where $\beth_{c}^{k+1}\subseteq\mathcal{S}_{c}^{k+1}$ is the set of participating servers for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$.
Case II-A. $\beth_{c}^{k+1}\cap(\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1})\neq\emptyset$, i.e., at least one server in $\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1}$ is chosen for assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$. Taking any server $m\in\beth_{c}^{k+1}\cap(\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1})$, we have
[…]
As a result, $\textsc{WF}_{k+1}-\textsc{WF}_{k}<\textsc{WF}_{k+1}<\textsc{OPT}$.
Case II-B. $\beth_{c}^{k+1}\cap(\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1})=\emptyset$. It follows that $\beth_{c}^{k+1}\subseteq\Omega_{c}^{k+1}$, which implies
$\beth_{c}^{k+1}=\Omega_{c}^{k+1}\cap\beth_{c}^{k+1}.$ (27)
We divide the servers in $\Omega_{c}^{k+1}$ into two disjoint sets:
$\Omega_{c,\leq}^{k+1}:=\left\{m\in\Omega_{c}^{k+1}\mid b_{m}^{c}(k)\leq\textsc{WF}_{k}\right\},$ (28)
$\Omega_{c,>}^{k+1}:=\left\{m\in\Omega_{c}^{k+1}\mid b_{m}^{c}(k)>\textsc{WF}_{k}\right\}.$ (29)
For each $m\in\Omega_{c,>}^{k+1}$, if server $m$ is a participating server in the assignment of any group $1,\ldots,k$ by WF, we have
[…]
giving a contradiction. Thus, server $m$ is not a participating server in the assignment of groups $1,\ldots,k$ by WF. As a result, $b_{m}^{c}(k)=b_{m}^{c}(0)$ holds for each $m\in\Omega_{c,>}^{k+1}$.
Case II-B1. $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}=\emptyset$. In this case, we have $\beth_{c}^{k+1}\subseteq\Omega_{c,\leq}^{k+1}$. In $\textsc{OPT}_{k+1}$, $\textsc{OPT}_{k+1}$ time slots of all the participating servers $\beth_{c}^{k+1}$ can accommodate all the tasks of $\mathcal{T}_{c}^{k+1}$. Thus,
$|\mathcal{T}_{c}^{k+1}|\leq\textsc{OPT}_{k+1}\cdot\sum_{m\in\beth_{c}^{k+1}}\mu_{m}^{c}.$ (31)
On the other hand, by the definition of WF, $(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)$ time slots of all the participating servers $\Omega_{c,\leq}^{k+1}$ are not adequate for all the tasks of $\mathcal{T}_{c}^{k+1}$. Hence,
$\left(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1\right)\cdot\sum_{m\in\Omega_{c,\leq}^{k+1}}\mu_{m}^{c}<|\mathcal{T}_{c}^{k+1}|.$ (32)
Combining (31) and (32) and noting that $\beth_{c}^{k+1}\subseteq\Omega_{c,\leq}^{k+1}$, we obtain $\textsc{WF}_{k+1}-\textsc{WF}_{k}-1<\textsc{OPT}_{k+1}$. By (21), we have $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$.
Case II-B2. $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}\neq\emptyset$. In this case, at least one server in $\Omega_{c,>}^{k+1}$ is a participating server for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$. Hence, all the servers in $\Omega_{c,\leq}^{k+1}$ must also be participating servers for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$. In other words, $\Omega_{c,\leq}^{k+1}\subseteq\beth_{c}^{k+1}$.
Figure 7: Visualization of Case II-B2.
In $\textsc{OPT}_{k+1}$, $\textsc{OPT}_{k+1}$ time slots of all the participating servers $\beth_{c}^{k+1}$ can accommodate all the tasks of $\mathcal{T}_{c}^{k+1}$ and the outstanding tasks of previous jobs queued at the servers in $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}$. Thus,
$|\mathcal{T}_{c}^{k+1}|+\sum_{m\in\Omega_{c,>}^{k+1}\cap\beth_{c}^{k+1}}b_{m}^{c}(0)\cdot\mu_{m}^{c}\leq\textsc{OPT}_{k+1}\cdot\sum_{m\in\beth_{c}^{k+1}}\mu_{m}^{c}.$ (33)
On the other hand, by the definition of WF, $(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)$ time slots of all the participating servers in $\Omega_{c}^{k+1}\cap\beth_{c}^{k+1}$ are not adequate for all the tasks of $\mathcal{T}_{c}^{k+1}$ and the backlogs beyond $\textsc{WF}_{k}$ (see the black box in Fig. 7 for an illustration). Hence,
$(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)\cdot\sum_{m\in\beth_{c}^{k+1}}\mu_{m}^{c}$
=(27)${=}(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)\cdot\sum_{m\in\Omega_{c}^{k+1}\cap\beth_{c}^{k+1}}\mu_{m}^{c}$
$\qquad<|\mathcal{T}_{c}^{k+1}|+\sum_{m\in\Omega_{c,>}^{k+1}\cap\beth_{c}^{k+1}}(b_{m}^{c}(0)-\textsc{WF}_{k})\cdot\mu_{m}^{c}.$ (34)
Note that the right side of (34) is no greater than the left side of (33). Combining (33) with (34), we obtain $\textsc{WF}_{k+1}-\textsc{WF}_{k}-1<\textsc{OPT}_{k+1}$. By (21), we have $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$.
With the above analysis for Case I and Case II, we have shown that (19) indeed holds. ∎
Theorems 1 and 2 together show that WF has a tight approximation ratio of $K_{c}$.
III-C Replica-Deletion
In this section, we introduce another task assignment heuristic called RD (replica-deletion).
III-C1 Algorithm Design
We illustrate RD with an example, as shown in Fig. 8. When a new job $c$ arrives, each task $r$ in $\mathcal{T}_{c}$ is initially replicated $|\mathcal{S}^{r}|$ times at all of its available servers. For instance, the blue task in Fig. 8 has three available servers: servers 1, 2, and 5, so it is replicated at these three servers in Fig. 8(i).
Figure 8: How RD works. (i)–(ix) demonstrate how to delete task replicas. In this example, $\mu_{m}^{c}\equiv 1$ for all the servers $m\in\mathcal{M}$.
After initialization, RD enters the deletion phase. To minimize the completion time of job $c$, RD removes redundant task replicas while keeping the estimated busy times of participating servers as balanced as possible. In each deletion iteration, RD first identifies the participating server(s) with the largest estimated busy time, referred to as the target server(s), denoted by $m^{\star}$. For example, in the first deletion iteration (Fig. 8(i)), the target server is server 2. RD then removes up to $\mu_{m^{\star}}^{c}$ task replicas from the target server to reduce its estimated busy time by one time slot. In each deletion iteration, RD removes task replicas with the largest numbers of copies.
In the example in Fig. 8, we assume $\mu_{m}^{c}\equiv 1$ for all the servers, so RD will remove at most one task replica from the target server. In Fig. 8(i), at server 2, the red task has two replicas (in servers 2 and 5), the blue task has three replicas, and the pink task has two replicas. Therefore, RD deletes the blue task replica from server 2. In the second deletion iteration (Fig. 8(ii)), the target servers are servers 1 and 2. In this case, the green task has the maximum number of replicas among all the tasks in these target servers. Thus, a green task replica is removed from server 1. In the third deletion iteration (Fig. 8(iii)), the target server is server 2, and both the red task and the pink task have two replicas. In this case, RD randomly removes one of them. Figs. 8(iv) to 8(vii) show the subsequent deletion iterations.
If there are multiple tasks with the same maximum number of replicas on distinct target servers, RD selects the server with the largest initial estimated busy time, i.e., the estimated busy time of servers before assigning any tasks of job $c$, as the target server. For example, in Fig. 9, the target servers are servers 1 and 5, and both the blue task and the green task have three replicas. In this case, the blue task replica on server 1 will be deleted since server 1 has a larger initial estimated busy time.
Figure 9: A target server with a larger initial estimated busy time has a task replica removed.
RD exits the deletion phase when all the tasks in the target server(s) have only one replica, indicating that the completion time of job $c$ cannot be reduced further. In Fig. 8(viii), the deletion phase terminates since all the tasks in target servers 1 and 5 have no other replicas. In the final phase, RD removes redundant task replicas from the remaining participating servers so that each task is processed by only one server. This removal procedure operates similarly to the deletion phase, identifying and removing replicas from the server with the largest estimated busy time among the remaining servers. For example, in Fig. 8(ix), RD deletes the replicas of the yellow and green tasks from server 4. This is done to keep the estimated busy times of all the servers as small and balanced as possible.
RD assigns tasks to servers from a global perspective because it scans all available servers when selecting a task replica to delete. In contrast, WF minimizes the completion time of the arrived job locally by dividing tasks into task groups and balancing only the available servers within each task group in each iteration. Intuitively, RD can outperform WF in most cases, but RD’s procedure is more complex to execute.
III-C2 Complexity Analysis
In the implementation of RD, we use a priority queue to maintain all servers, prioritizing by their estimated busy times (and break ties by their initial estimated busy times). Additionally, for each server $m$, we use a priority queue to maintain the assigned tasks, prioritizing by their numbers of replicas. In the worst case, each task is replicated on all servers, requiring $\mathcal{O}(|\mathcal{T}_{c}|\cdot M)$ deletions. Moreover, each deletion affects the priority queues of the servers containing the replicas of the deleted task. Thus, each deletion leads to updating the priority queue of up to $M$ servers, resulting in a complexity of $\mathcal{O}(M\cdot\log|\mathcal{T}_{c}|)$. Therefore, the total time complexity of RD is $\mathcal{O}(M^{2}\cdot|\mathcal{T}_{c}|\cdot\log|\mathcal{T}_{c}|)$. We experimentally compare the performance of RD with WF and other algorithms in Sec. V, and leave its theoretical analysis as future work.
IV Algorithms with Job Reordering
In this section, we consider an extended scenario in which the outstanding job tasks queued in each server can be prioritized and reordered. By adjusting the execution order of outstanding jobs and reassigning their remaining tasks, their average completion time can be further reduced.
Inspired by [4] and [7], we use a similar approach to derive the job execution order, which emulates a shortest-remaining-time-first policy for distributed job executions. When a new job $c$ arrives, job reordering is triggered. We denote by $\mathcal{O}_{c}$ the set of outstanding jobs after job $c$’s arrival and denote by $\mathcal{Q}_{c}$ the reordered list of outstanding jobs. Our goal is to retrieve jobs from $\mathcal{O}_{c}$ and insert them into $\mathcal{Q}_{c}$ in a specific sequence until $|\mathcal{Q}_{c}|=|\mathcal{O}_{c}|$.
Assume that there are $p$ sorted outstanding jobs in $\mathcal{Q}_{c}$. For each job $c^{\prime}\in\mathcal{O}_{c}\backslash\mathcal{Q}_{c}$, we estimate the remaining time to complete job $c^{\prime}$ as $\Phi_{c^{\prime}}$, assuming that job $c^{\prime}$ is the $(p+1)$-th job in the new order. Here, $\Phi_{c^{\prime}}$ is derived by WF with the estimated busy times of servers, given the assignment of $p$ jobs in $\mathcal{Q}_{c}$ and the number of unprocessed tasks of job $c^{\prime}$. We denote by $l$ the index of the job with the minimal $\Phi_{l}$ among all the outstanding jobs explored. When exploring each job $c^{\prime}\in\mathcal{O}_{c}\backslash\mathcal{Q}_{c}$, we first compute a lower bound $\Phi_{c^{\prime}}^{-}$ on the completion time of job $c^{\prime}$, as indicated by (8) and (9) in Sec. III-A2. If $\Phi_{c^{\prime}}^{-}\geq\Phi_{l}$, the exploration of job $c^{\prime}$ terminates. Otherwise, we calculate $\Phi_{c^{\prime}}$ by WF and update the best-explored job $l$ if $\Phi_{c^{\prime}}<\Phi_{l}$. The reason is that if $\Phi_{c^{\prime}}^{-}>\Phi_{l}$, letting job $c^{\prime}$ be the $(p+1)$-th job can never lead to a lower average job completion time. This technique, which significantly reduces the computation overhead on finding the $(p+1)$-th job in the new order, is named early-exit.
1 while a new job $c$ arrives do
2       $\mathcal{O}_{t}\leftarrow\mathcal{O}_{t}\cup\{c\}$
3       $\mathcal{Q}_{t}\leftarrow\emptyset$
4       Initialize the estimated busy times of servers: $\forall m\in\mathcal{M}$, $b_{m}^{c}\leftarrow 0$
5       while $|\mathcal{Q}_{c}|<|\mathcal{O}_{c}|$ do
6             $l\leftarrow\texttt{nullptr}$
7             for each job $c^{\prime}\in\mathcal{O}_{t}\backslash\mathcal{Q}_{t}$ do
8                   Get the unprocessed tasks of job $c^{\prime}$, and calculate $\Phi_{c^{\prime}}^{-}$ by (8) and (9)
9                   if $l==\texttt{nullptr}$ or $\Phi_{c^{\prime}}^{-}<\Phi_{l}$ then
10                         Get the unprocessed tasks of job $c^{\prime}$, and solve the task assignment problem $\mathcal{P}$ by WF to get $\Phi_{c^{\prime}}$
11                         if $l==\texttt{nullptr}$ or $\Phi_{c^{\prime}}<\Phi_{l}$ then
12                               Derive the assignment of unprocessed tasks by WF
13                               $l\leftarrow c^{\prime}$
14
15                   else
                        break // early-exit
16
17
18            $\mathcal{Q}_{c}\leftarrow\mathcal{Q}_{c}\cup\{l\}$
19             For each $m\in\mathcal{M}$, update the estimated busy time of server $m$
20
Algorithm 3 OCWF-ACC
We name the algorithm as OCWF-ACC (accelerated order-conscious WF). Algorithm 3 presents the details of OCWF-ACC. Note that WF can be replaced by other task assignment algorithms.
V Experimental Results
We conduct extensive simulations to verify the performance and efficiency of the proposed algorithms. We first present the simulation settings. Then, we show and discuss the results.
V-A Experimental Setup
Figure 10: Average job completion time, computation overheads, and CDF of job completion times under 25% system utilization for different algorithms.
Figure 11: Average job completion time, computation overheads, and CDF of job completion times under 50% system utilization for different algorithms.
Figure 12: Average job completion time, computation overheads, and CDF of job completion times under 75% system utilization for different algorithms.
Job Traces. We use the Alibaba cluster trace program [8] to drive the simulation. We extract a segment from the dataset batch_task.csv in cluster-trace-v2017 that contains 250 jobs. There are a total of 113,653 tasks in these jobs. We derive job arrivals and task durations from the timestamps of the recorded task events. We scale the inter-arrival times of the jobs to simulate different levels of system utilization, ranging from 25% to 75%. The default number of servers is 100. For each server, its profiled processing capacity for each job is randomly generated between 3 and 5 by default.
Available Servers. We consider each entry of a job in the trace file as one task group of the job. Each entry is a task event, which records the creation timestamp, job_id, number of task instances, etc. On average, there are 5.52 task groups in a job. The data inputs to task groups are assumed to be distributed among the servers according to a Zipf distribution. Specifically, for each task group, we first randomly generate a permutation of all servers. Then, the task group is associated with the $i$-th server in the permutation with a probability proportional to $\frac{1}{i^{\alpha}}$, where $\alpha\in[0,2]$ is the Zipf skew parameter ($\alpha=0$ means a uniform distribution). If the associated server of the task group is server $m$, then servers $m,m+1,\dots,m+p-1$ are chosen to be its available servers. Here, $p$ is randomly generated between 8 and 12 by default.
Algorithms. We implement six algorithms with DOcplex [^3] https://ibmdecisionoptimization.github.io/docplex-doc/index.html: NLIP, OBTA, WF, RD, OCWF, and OCWF-ACC. NLIP differs from OBTA in that it solves the non-linear program $\mathcal{P}$ for each job directly, without narrowing the search space of $\Phi_{c}$ and dividing it into subranges. For NLIP, OBTA, WF, and RD, we use the default FIFO policy for ordering job execution. OCWF is the non-accelerated version of OCWF-ACC, where the early-exit technique is not employed in the reordering of outstanding jobs (essentially the same as the implementation of the SWAG [4] and ATA-Greedy [7] algorithms). The code is available for tests in other settings or large-scale validations. [^4] https://github.com/hliangzhao/taos
Metrics. We use the average job completion time of all jobs to measure performance and the computation overhead of each algorithm to measure efficiency.
V-B Experimental Results
We compare the performance and efficiency of the algorithms by varying the Zipf parameter $\alpha$ from 0 to 2 and the system utilization from 25% to 75%. Figs. 10-12 present the results. In these figures, the right y-axis in the first subfigure indicates the average job completion time, while the left y-axis shows the log-scaled average computation overhead per job arrival. The other four subfigures visualize the cumulative distribution of job completion times for different $\alpha$ values. The key findings are summarized as follows:
Compared with NLIP, OBTA achieves a significant improvement in efficiency. From Figs. 10-12, we observe that OBTA and NLIP have fairly close performance of job completion time since both are theoretically optimal for balanced task assignment. By narrowing the search space and dividing it into subranges, OBTA reduces the computation overhead by nearly half compared to NLIP. This verifies our contribution in Sec. III-B.
WF closely approximates balanced task assignment with extremely low overhead. WF performs close to OBTA and NLIP at almost all percentiles of job completion times, and its computation overhead is two orders of magnitude lower than OBTA’s. The ultra-low overhead offers a significant advantage and facilitates job reordering.
RD’s performance positions it between WF and OBTA. Our results show that RD generally outperforms WF. This observation is detailed in Table I, which presents the performance of each algorithm when $\alpha=2$ and system utilization is at 75%, as shown in the first figure of Fig. 13. However, it is noteworthy that RD has a higher computation overhead than WF, which aligns with our complexity analysis. The time complexity of RD is quantified as $\mathcal{O}(M^{2}\cdot n\log n)$, which is higher than WF’s $\mathcal{O}(KM\cdot\log n)$, where $M$ is the number of servers, $n$ is the number of tasks, and $K$ is the number of task groups in a job.
TABLE I: Average job completion time vs. #available servers.
Algorithms Number of available servers
$4$ $6$ $8$ $10$ $12$ Aver.
OBTA $12551$ $5814$ $4217$ $3609$ $3160$ $5870$
NLIP $12468$ $5924$ $4394$ $3547$ $3164$ $5899$
WF $12872$ $6125$ $4408$ $3617$ $3188$ $6042$
RD $12857$ $5987$ $4261$ $3595$ $3150$ $5970$
OCWF $1699$ $958$ $765$ $712$ $657$ $958$
OCWF-ACC $1699$ $958$ $765$ $712$ $657$ $958$
OCWF-ACC and OCWF are robust to skewness in data availability. As shown in Figs. 10-12, when the skew parameter $\alpha$ changes from 0 to 2, the average job completion time achieved by FIFO-based algorithms (NLIP, OBTA, WF, and RD) shows a clear upward trend. In contrast, OCWF and OCWF-ACC do not exhibit a significant increase in average job completion time. This demonstrates that reordering and reallocation are crucial for handling skewness in data availability.
OCWF-ACC significantly accelerates OCWF. The computation overhead of OCWF-ACC is only about half of OCWF’s overhead. This verifies the advantage of the early-exit technique.
Figure 13: Average job completion time, and CDF of job completion times when $\alpha=2$ under 75% system utilization for different numbers of available servers.
Figure 14: Average job completion time, and CDF of job completion times when $\alpha=2$ under 75% system utilization for different processing capacities.
Figs. 10-12 show that job completion times increase with system utilization. The performance trends discussed above are consistently observed across different levels of system utilization.
Figs. 13-14 show the job completion times for varying numbers of available servers and profiled processing capacities, respectively. These experiments are conducted with $\alpha=2$ and a system utilization of 75%, as we are particularly interested in the performance when resource contention is high. From Fig. 13, we observe that a larger number of available servers decreases job completion times. More available servers provide greater flexibility in task assignment, allowing tasks to be distributed in a manner that better balances the load among servers and reduces job completion times. Fig. 14 demonstrates that increasing processing capacity also reduces job completion times. This is expected, as higher processing capacities enable more tasks to run in parallel, thereby decreasing job completion times. Notably, the relative performance of the algorithms remains largely unchanged across different numbers of available servers and processing capacities.
VI Related Work
Job scheduling is extensively studied from both theoretical [9, 10, 11, 12, 13, 14, 15, 16] and system-level perspectives [17, 18, 19, 20, 21, 22]. Theoretical works often formulate job completion time minimization problems as combinatorial, constrained optimization problems, solving them with various approaches, especially approximate algorithms. Data locality refers to executing a job close to where the requested data chunk resides, instead of moving data to computation. Originating from Map-Reduce [3], many algorithms have been proposed in both homogeneous and heterogeneous settings [23, 24, 25, 6]. Generally, these works aim to balance data locality with scenario-oriented performance indicators such as load balancing [26], throughput [27], delay [25, 28], fairness [5], and job completion time [29, 4, 30].
Preserving data locality, some works have studied the scheduling of distributed job executions [4, 31, 5, 6, 7]. In these scenarios, jobs arrive online with multiple tasks, and each task requires a data chunk that may be replicated across different servers. To minimize job completion times, Hung et al. [4] proposed a workload-aware greedy scheduling algorithm called SWAG, which orders job executions by estimating the completion times of outstanding jobs. SWAG is the first algorithm to consider the order of outstanding jobs in distributed job executions. Guan et al. [7] significantly improved SWAG with a BTAaJ algorithm, which assigns tasks of a newly arrived job based on a well-constructed flow network. BTAaJ groups the tasks of a job by their available servers and achieves load balancing by minimizing a control parameter $C$. They also proposed an ATA-Greedy algorithm to improve scheduling performance by reordering outstanding jobs using a similar approach to [4], but with the control parameter $C$ to estimate completion times. Our work presents significant improvements over [4] and [7].
The scheduling of distributed job executions has also been considered from other perspectives, such as fairness [5, 31] and data movement [6]. Guan et al. [5] extended conventional max-min fairness on a single machine to distributed job execution over multiple servers, proposing a policy called Aggregate Max-min Fairness, which is Pareto-efficient, envy-free, and strategy-proof. Beaumont et al. [6] studied a problem called Comm-Oriented, which enforces the optimality of makespan and minimizes data movement. However, this work focused on a single job only and is not applicable to our problem.
VII Conclusion
In this paper, we explored the task assignment problem in the context of online distributed job executions, with data locality preserved. Our contributions began with formulating the task assignment problem as a non-linear integer program. We developed the OBTA algorithm, known for its efficiency in reducing computational overhead by narrowing the search space of potential solutions. Additionally, we extended the WF heuristic, providing an in-depth, nontrivial tight analysis of its approximation factor, and introduced the innovative RD algorithm, which demonstrates enhanced performance. Our investigation into job reordering, particularly with the implementation of the novel early-exit technique, effectively optimized average job completion times while maintaining low computational overheads. The effectiveness of all proposed algorithms was validated through extensive trace-driven simulations. Both OBTA and the early-exit technique showed substantial reductions in computational overhead. Meanwhile, WF emerged as a practical, lower-overhead alternative for large-scale scenarios. Despite its greater computational intensity compared to WF, RD exhibited superior performance of job completion time. Looking ahead, integrating these algorithms into diverse computational settings and adapting them to the evolving dynamics of distributed job executions hold potential for further enhancing their applicability.
Acknowledgments
This work was supported in part by the Singapore Ministry of Education under Academic Research Fund Tier 2 Award MOE-T2EP20121-0005, the National Natural Science Foundation of China under Grants 62125206 and U20A20173, the National Key Research and Development Program of China under Grant 2022YFB4500100, and the Key Research Project of Zhejiang Province under Grant 2022C01145.