### v4

# Data-Locality-Aware Task Assignment and Scheduling for Distributed Job Executions

Abstract
This paper addresses the data-locality-aware task assignment and scheduling problem for distributed job executions. Our goal is to minimize job completion times without prior knowledge of future job arrivals. We propose an Optimal Balanced Task Assignment algorithm (OBTA), which achieves minimal job completion times while significantly reducing computational overhead through efficient narrowing of the solution search space. To balance performance and efficiency, we extend the approximate Water-Filling (WF) algorithm, providing a rigorous proof that its approximation factor equals the number of task groups in a job. We also introduce a novel heuristic, Replica-Deletion (RD), which outperforms WF by leveraging global optimization techniques. To further enhance scheduling efficiency, we incorporate job ordering strategies based on a shortest-estimated-time-first policy, reducing average job completion times across workloads. Extensive trace-driven evaluations validate the effectiveness and scalability of the proposed algorithms.
Index Terms:
task assignment, job scheduling, distributed job execution, approximate analysis, trace-driven evaluation.
I Introduction
Scheduling is a fundamental component of modern service-oriented computing, enabling distributed systems to efficiently execute diverse workloads and deliver scalable services [1, 2, 3, 4]. In cloud-based environments, jobs typically comprise multiple tasks that require access to specific data chunks, which are often replicated across geographically distributed servers. Efficient data-locality-aware scheduling, which assigns tasks to servers hosting the required data chunks, is critical for reducing data transfer overhead, enhancing task execution efficiency, and ensuring compliance with stringent service-level agreements (SLAs) [5]. The design of online scheduling algorithms for such multi-task jobs introduces two central challenges: task assignment and job ordering. Task assignment entails mapping tasks to appropriate servers based on data locality, often modeled as a semi-matching problem on a bipartite graph with tasks and servers as nodes. Job ordering involves determining the execution sequence of pending tasks to optimize resource utilization and minimize job completion times, a key requirement in dynamic and resource-intensive service environments.
Existing research has made progress on these challenges, yet often under assumptions that limit applicability in real-world service systems [6, 7, 8, 9, 3, 4]. For instance, Hung et al. [6] proposed a workload-aware greedy scheduling strategy (SWAG) that prioritizes jobs based on estimated completion times. While effective in specific scenarios, this approach neglects data replication, a prevalent feature in distributed service systems. Similarly, Guan et al. [7] introduced an offline fairness-aware task assignment algorithm to ensure max-min fairness across servers, but their method lacks the adaptability needed for online and dynamic service platforms. Beaumont et al. [8] explored scheduling with data replication but constrained their analysis to single-job scenarios and homogeneous environments, which oversimplifies the heterogeneity and complexity inherent in service-oriented infrastructures. Guan et al. [9] tackled an online scheduling problem for distributed job executions, leveraging maximum flow-based techniques for task assignment. However, their work offered limited performance analysis and did not thoroughly examine job ordering strategies essential for dynamic service environments.
In this paper, we focus on developing data-locality-aware task assignment and scheduling algorithms with theoretical guarantees tailored for online distributed service environments. Our approach addresses two fundamental scenarios: (i) jobs are executed in a FIFO (first-in-first-out) manner, and (ii) jobs can be prioritized and reordered to optimize performance. By addressing these scenarios, we aim to improve the efficiency of task scheduling in service-oriented systems, tackling the dual challenges of data locality and dynamic workload management. The main contributions of this work are summarized as follows:
  • We formulate the task assignment problem as a bipartite graph-based non-linear program and propose an Optimal Balanced Task Assignment algorithm (OBTA), which significantly reduces the search space and provides efficient solutions for heterogeneous service environments.
  • We extend the approximate water-filling (WF) algorithm, originally proposed for homogeneous settings in [9], to heterogeneous settings. We provide a nontrivial proof that WF achieves a $K$-approximation, where $K$ represents the number of task groups in a job.
  • We propose a novel heuristic, replica-deletion (RD), for task assignment. RD generally produces better task assignments than WF.
  • We extend WF to support job reordering and introduce an accelerated order-conscious WF (OCWF-ACC) algorithm, which incorporates an early-exit technique to improve computational efficiency.
  • We validate the proposed algorithms through extensive trace-driven evaluations. OBTA demonstrates significant reductions in average job completion times, while WF provides a low-overhead yet effective approximation. RD enhances overall performance further. Moreover, OCWF-ACC achieves substantial reductions in computation overhead compared to job-reordered task assignment algorithms without early-exit techniques.
The remainder of this paper is organized as follows: Sec. II defines the problem formulation. Sec. III presents task assignment algorithms designed for FIFO scheduling. Sec. IV extends these algorithms to support job reordering integration. Sec. V provides experimental results based on real-world job traces, demonstrating the effectiveness of the proposed methods. Sec. VI reviews related work, highlighting the context and contributions of this study. Finally, Sec. VII concludes the paper and discusses potential future research directions.
II Problem Formulation
We consider a distributed computing system comprising $M$ servers, indexed by $m$ for $m\in\mathcal{M}:=\{1,...,M\}$. The system stores a collection of data items, such as key-value pairs, objects, or files, partitioned into equally-sized data chunks. Each server holds a subset of these data chunks, and each chunk can be replicated across multiple servers. For each server $m\in\mathcal{M}$, let $\mathcal{D}_{m}$ denote the set of data chunks stored on it, where $\mathcal{D}_{m}\cap\mathcal{D}_{m’}$ can be non-empty for $m\neq m’$ ’. Key notations and acronyms used in this paper are summarized in Table I and Table II, respectively.
TABLE I: Summary of key notations.
Notation Description
$\mathcal{M}$ The set of servers
$m$ Index of servers in $\mathcal{M}$
$\mathcal{D}_{m}$ The set of data chunks in server $m\in\mathcal{M}$
$c$ Index of arriving jobs
$\mathcal{T}_{c}$ The set of tasks in job $c$
$r$ Index of tasks
$\mathcal{S}^{r}$ The available servers (due to data locality) of task $r$
$\mu_{m}^{c}$ Processing capacity of server $m$ for job $c$
$\Phi_{c}$ The estimated completion time of job $c$
$\Phi_{c}^{+}(\Phi_{c}^{-})$ The upper (lower) bound of $\Phi_{c}$ derived by OBTA
$b_{m}^{c}$ The busy time of server $m$ at the arrival of job $c$
$K_{c}$ The number of task groups in job $c$
$k$ Index of task groups
$\mathcal{T}_{c}^{k}$ The set of tasks in the $k$-th group of job $c$
$\mathcal{S}_{c}^{k}$ The set of available servers of the task group $\mathcal{T}_{c}^{k}$
ALG An algorithm ALG (and its objective value)
$\alpha_{\textsc{ALG}}$ The approximation ratio of the algorithm ALG
$I$ A job arrival instance
$\mathcal{O}_{c}$ The set of outstanding jobs at the arrival of job $c$
TABLE II: Summary of acronyms of related algorithms. OBTA-N and OBTA-P are introduced in Sec. VI.
Acronym Description
SWAG workload-aware greedy scheduling strategy
OBTA optimal balanced task assignment
NLIP non-linear integer programming
WF water-filling
RD replica deletion
OCWF order-conscious water-filling
OCWF-ACC accelerated order-conscious water-filling
LIP linearized integer programming
OBTA-N OBTA with only solution space narrowing
OBTA-P OBTA with only piecewise linearization
Jobs arrive dynamically, one at a time, for execution. Each job consists of multiple independent tasks, where each task processes a specific data partition. A task requires exactly one data chunk as input, and all tasks within a job are assumed to have identical computational demands. Let $c\in\mathbb{N}^{+}$ denote the $c$-th job (also referred to as job $c$) in chronological order. Further, let $\mathcal{T}_{c}$ represent the set of tasks in job $c$, and let $d_{r}$ denote the data chunk required by a task $r\in\mathcal{T}_{c}$. For each task $r\in\mathcal{T}_{c}$, the set of available servers capable of processing the task is defined as:
$\mathcal{S}^{r}:=\big{\{}m\in\mathcal{M}\mid d_{r}\in\mathcal{D}_{m}\big{\}}.$ (1)
To maintain data locality, each task must be assigned to one of its available servers for execution. Each server $m$ maintains a queue $q_{m}$ to store outstanding tasks, with no predefined limit on queue size. For analysis purposes, the system’s time is divided into discrete, identical time slots.
Each server $m$ has a task processing capacity characterized by $\mu_{m}^{c}$, representing the number of tasks from job $c$ that server $m$ can process within a single time slot. The initial busy time of each server $m$ just before job $c$’s arrival, denoted by $b_{m}^{c}$, is the number of time slots required to process all tasks currently queued at server $m$. This busy time can be estimated using:
$b_{m}^{c}=\sum_{h=1}^{c-1}\left\lceil\frac{o_{m}^{h}}{\mu_{m}^{h}}\right\rceil,$ (2)
where $o_{m}^{h}\leq|\mathcal{T}_{h}|$ is the number of tasks from job $h$ that remain pending in $q_{m}$ when job $c$ arrives.
The objective is to design a scheduling strategy that assigns tasks to servers and determines execution priorities to minimize the completion time of each job. The completion time of a job is defined as the duration between its arrival and the completion of its last processed task, equivalent to minimizing the finish time of the last task in the job. We investigate the problem under two distinct scenarios:
  • FIFO queues. Outstanding tasks are processed in the order of their arrival (first-in-first-out). The challenge is to assign tasks of each arriving job to available servers while balancing server workloads to minimize job completion times.
  • Prioritized reordering. Outstanding tasks can be dynamically reordered in each server’s queue. This scenario focuses on prioritizing tasks to further reduce job completion times, introducing additional complexity as reordering decisions influence overall system performance.
We do not impose any specific assumptions about the timing or pattern of job arrivals.
III Algorithms for FIFO Queues
III-A Matching-based Non-Linear Programming
III-A1 Programming in a Bipartite Graph
When a new job $c$ arrives, the task assignment problem involves allocating all tasks in $\mathcal{T}_{c}$ to appropriate servers while ensuring data locality and minimizing the job completion time.
To facilitate this, we divide the tasks into task groups, where each task group contains tasks sharing the same set of available servers. Let $K_{c}$ denote the number of task groups for job $c$. Further, for each $k\in\mathcal{K}_{c}:=\{1,\ldots,K_{c}\}$, we denote by $\mathcal{T}_{c}^{k}$ and $\mathcal{S}_{c}^{k}$ the set of tasks in the $k$-th group (also referred to as group $k$) and the set of available servers for tasks in the $k$-th group, respectively. By definition, we have
$\mathcal{T}_{c}^{k}=\left\{r\in\mathcal{T}_{c}\mid\mathcal{S}^{r}=\mathcal{S}_{c}^{k}\right\}.$ (3)
[…]
$\mathcal{P}:\qquad\qquad\qquad\min_{\Phi_{c},n_{m}^{k}\in\mathbb{N}^{+}}\Phi_{c}$
s.t. $\left\{\begin{array}[]{ll}\sum_{k\in\mathcal{K}_{c}}n_{m}^{k}\leq\max\left\{\Phi_{c}-b_{m}^{c},0\right\}&\forall m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k},\\ \sum_{m\in\mathcal{S}_{c}^{k}}n_{m}^{k}\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|&\forall k\in\mathcal{K}_{c}.\end{array}\right.$ (6)
[…]
III-A2 Narrowing the Search of $\Phi_{c}$
Commercial solvers can be computationally expensive when solving non-linear integer programs, especially in large-scale scenarios. To mitigate this overhead, we narrow the search space of $\Phi_{c}$ by leveraging the inherent characteristics of the problem.
After the arrival of job $c$, the total number of time slots required to process both the backlogged tasks and the newly assigned tasks from job $c$ is bounded above by:
$\Phi_{c}^{+}:=\max_{m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}}\left\{\left\lceil\frac{\sum_{k\in\mathcal{K}_{c}:m\in\mathcal{S}_{c}^{k}}|\mathcal{T}_{c}^{k}|}{\mu_{m}^{c}}\right\rceil+b_{m}^{c}\right\}.$ (7)
This upper bound, $\Phi_{c}^{+}$, is determined under the assumption that all tasks in $\mathcal{T}_{c}$ are assigned to a single available server $m$. The corresponding lower bound of $\Phi_{c}$ is defined as:
$\Phi_{c}^{-}:=\max_{k\in\mathcal{K}_{c}}x_{k},$ (8)
where $x_{k}$ is the minimum integer satisfying:
$\sum_{m\in\mathcal{S}_{c}^{k}}\left(\max\left\{x_{k}-b_{m}^{c},0\right\}\right)\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|.$ (9)
Here, $x_{k}$ represents the minimal number of time slots required to process all tasks in $\mathcal{T}_{c}^{k}$, assuming that task group $k$ is the only group being processed for job $c$. By combining these bounds, we reduce the search space of $\Phi_{c}$ from $\mathbb{N}^{+}$ to the integer interval $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$. This narrowing significantly reduces the computational complexity of solving the task assignment problem, making it more tractable for large-scale systems.
III-A3 The Matching-based Algorithm
Figure 1: Divide $\big{[}\Phi_{c}^{-},\Phi_{c}^{+}\big{]}$ into disjoint subranges.
By introducing $\Phi_{c}^{-}$ and $\Phi_{c}^{+}$, the search space of $\Phi_{c}$ is significantly narrowed. However, the problem $\mathcal{P}$ remains non-linear due to the piecewise nature of the first constraint in (6). To address this, we divide the interval $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$ into multiple disjoint sub-intervals based on the initial estimated busy times of the available servers. This division is illustrated in Fig. 1. Specifically, the initial estimated busy times $b_{m}^{c}$ of all servers $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ are sorted in ascending order, and the sorted values are denoted as $b_{1}’,b_{2}’,\ldots$ ’ , …. We then identify indices $i$ and $j$, where $i$ is the smallest index such that $b_{i}’\geq\Phi_{c}^{-}$, and $j$ is the largest index such that $b_{j}’\leq\Phi_{c}^{+}$. With this setup, the search space of $\Phi_{c}$ is divided into sub-intervals: $\left[\Phi_{c}^{-},b_{i}’\right)$ ’ ), $\left[b_{i}’,b_{i+1}’\right)$ ’ ), …, $\left[b_{j}’,\Phi_{c}^{+}\right]$. Within each sub-interval, the piecewise nature of the first constraint in (6) is resolved, transforming $\mathcal{P}$ into a linear integer program. These sub-problems are then solved sequentially. If $\mathcal{P}$ is solvable within a particular sub-interval, the corresponding solution is guaranteed to be optimal, as the remaining sub-intervals cannot yield a smaller $\Phi_{c}$.
Using this approach, we propose the Optimal Balanced Task Assignment (OBTA) algorithm. When the profiled values $\{\mu_{m}^{c}\}_{m,c}$ are accurate, OBTA produces optimal solutions by computing the exact values of $n_{m}^{k}$ for task assignment. The pseudocode for OBTA is provided in Algorithm 1.
Input: Online arriving jobs $1,2,...,c,...$, server capacities $\{\mu_{m}^{c}\}_{m,c}$
Output: Assignment solution for each task of each arriving job
1 while a new job $c$ arrives do
2 foreach $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
3      Estimate $b_{m}^{c}$ using (2)
4  Solve $\mathcal{P}$ as described in Sec. III-A3, obtaining $\{n_{m}^{k}\}_{m,k}$
5 foreach $k\in\mathcal{K}_{c}$ do
6    foreach $m\in\mathcal{S}_{c}^{k}$ do
7       if server $m$ is not the last server in $\mathcal{S}_{c}^{k}$ then
[…]
9       else
[…]
11
12
13
Algorithm 1 OBTA
III-B Water-Filling
Solving a group of integer programs can still be time-consuming for large-scale services environments. To improve efficiency, we propose an alternative task assignment algorithm called Water-Filling (WF), which is extended from [9]. WF assigns tasks to jobs sequentially, one group at a time, using an approximate approach to balance computational efficiency with performance.
III-B1 Algorithm Design
WF dynamically estimates server workloads as tasks are assigned. Let $b_{m}^{c}(k)$ denote the estimated busy time of server $m$ after assigning tasks from group $k$. Initially, for each server $m\in\mathcal{M}$, the busy time is:
$b_{m}^{c}(0):=b_{m}^{c},$ (10)
and it increases monotonically as task groups are assigned:
$b_{m}^{c}(0)\leq b_{m}^{c}(1)\leq\cdots\leq b_{m}^{c}\left(K_{c}\right).$ (11)
For each task group $k$, WF determines the minimum time $\xi_{k}$ required to process all tasks in the group across its available servers. Specifically, $\xi_{k}$ is the smallest integer satisfying:
$\sum_{m\in\mathcal{S}_{c}^{k}}\left(\max\left\{\xi_{k}-b_{m}^{c}(k-1),0\right\}\right)\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|,$ (12)
where $\mathcal{S}_{c}^{k}$ is the set of servers available for group $k$, and $b_{m}^{c}(k-1)$ represents the server’s busy time after the previous task group assignment. The calculation of $\xi_{k}$ is analogous to $x_{k}$ in (9), representing the estimated time needed to process all tasks in the group. A server $m\in\mathcal{S}_{c}^{k}$ is considered a participating server if $\xi_{k}-b_{m}^{c}(k-1)>0$. We then allocate $(\xi_{k}-b_{m}^{c}(k-1))\cdot\mu_{m}^{c}$ tasks in group $k$ to each participating server $m$ (or all the remaining tasks in group $k$ if server $m$ is the last participating server). After the assignment, the busy time of each server is updated as:
$b_{m}^{c}(k)=\max\left\{b_{m}^{c}(k-1),\xi_{k}\right\}.$ (13)
This process, repeated for each task group, simulates a water-filling mechanism where the participating servers’ workloads are gradually equalized. Fig. 2 illustrates the water-filling concept for task assignment.
Figure 2: Assign tasks in $\mathcal{T}_{c}^{k}$ to their available servers in a water-filling manner. In (i), only a subset of $\mathcal{S}_{c}^{k}$, including servers $1,...,m^{*}$, participate in the assignment of $\mathcal{T}_{c}^{k}$; In (ii), every server in $\mathcal{S}_{c}^{k}$ participates in the assignment of $\mathcal{T}_{c}^{k}$.
The pseudocode of WF is presented in Algorithm 2. Its computational complexity is $\mathcal{O}\left(\sum_{k\in\mathcal{K}_{c}}|\mathcal{S}_{c}^{k}|\log(|\mathcal{S}_{c}^{k}|)\right)$, as $\xi_{k}$ can be calculated via binary search with $\mathcal{O}(\log(|\mathcal{S}_{c}^{k}|))$ iterations. This complexity can also be expressed as $\mathcal{O}(K_{c}\cdot M\cdot\log M)$.
Input: Online arriving jobs $1,2,...,c,...$, server capacities $\{\mu_{m}^{c}\}_{m,c}$
Output: Assignment solution for each task of each arriving job
1 while a new job $c$ arrives do
2 for each $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
3      Estimate $b_{m}^{c}$ by (2)
4    $b_{m}^{c}(0)\leftarrow b_{m}^{c}$
5 for each $k\in\mathcal{K}_{c}$ do
6      Calculate $\xi_{k}$ as detailed in Sec. III-B1
7    for each $m\in\mathcal{S}_{c}^{k}$ do
8       if $b_{m}^{c}(k-1)\geq\xi_{k}$ then
9          continue
10       if server $m$ is not the last server in $\mathcal{S}_{c}^{k}$ then
[…]
12       else
[…]
14        Update $b_{m}^{c}(k)$ by (13)
15
16
Algorithm 2 WF
III-B2 Approximation Analysis
In this section, we analyze the performance of WF. The approximation factor of WF, compared to the optimal algorithm, is defined as:
$\alpha_{\textsc{WF}}:=\max_{I}\frac{\text{WF}(I)}{\text{OPT}(I)},$ (14)
where $I:=I\big{(}c,\{b_{m}^{c}\}_{m}\big{)}$ represents the arrival instance of a new job $c$ and the initial estimated busy times of all servers prior to job $c$’s arrival. Here, $\text{OPT}(I)$ and $\text{WF}(I)$ denote the maximum estimated busy times of participating servers after assigning job $c$’s tasks using the optimal algorithm and WF, respectively.
Figure 3: Visualization of the task assignment by WF for a constructed instance.
Theorem 1.
For each newly arrived job $c$, the approximation factor of WF is at least $K_{c}$.
Proof.
The WF-to-optimal ratio is 1 under favorable conditions, such as when $\mathcal{T}_{c}$ consists of only one task group or when $\mathcal{S}_{c}^{k}\cap\mathcal{S}_{c}^{k’}=\emptyset$ for all $k\neq k’$ ’. However, to demonstrate the worst-case ratio, we construct a specific instance $I$, illustrated in Fig. 3, where the available servers of different task groups overlap.
For simplicity, let $\mu_{m}^{c}=1$ and $b_{m}^{c}=0$ for all servers $m$. For each task group $k$, the number of its available servers is defined as:
$|\mathcal{S}_{c}^{k}|=\sum_{k^{\prime}=1}^{K_{c}-k+1}x^{k^{\prime}},$ (15)
where $x\geq 2$ is an integer. Furthermore, the available servers follow a hierarchical structure:
$\mathcal{S}_{c}^{1}\supset\mathcal{S}_{c}^{2}\supset\cdots\supset\mathcal{S}_{c}^{K_{c}},$ (16)
meaning the servers available for a higher-indexed group are a subset of those available for a lower-indexed group. Each task group $k$ contains
$|\mathcal{T}_{c}^{k}|=x\cdot|\mathcal{S}_{c}^{k}|$
tasks. In WF, tasks in each group are assigned sequentially, and as shown in Fig. 3, $x$ time slots are required to process all tasks in each group.
Figure 4: Visualization of the task assignment by OPT for the constructed instance.
In the optimal assignment (OPT), illustrated in Fig. 4, tasks from group $k$ ($k<K_{c}$) are assigned to a subset of its servers $\mathcal{S}_{c}^{k}\setminus\mathcal{S}_{c}^{k+1}$. The number of time slots required to process all tasks in group $k$ is:
$\Bigg{\lceil}\frac{x\cdot\sum_{i=1}^{K_{c}-k+1}x^{i}}{x^{K_{c}-k+1}}\Bigg{\rceil}=x+2.$ (17)
For group $K_{c}$, the number of time slots required remains $x$. Therefore, the total completion time for job $c$ using OPT is:
$\textsc{OPT}(I)=\max\{x+2,x\}=x+2.$ (18)
The approximation factor for this instance is:
$\frac{\textsc{WF}(I)}{\textsc{OPT}(I)}=\frac{K_{c}\cdot x}{x+2},$ (19)
As $x\to\infty$, the ratio approaches $K_{c}$, demonstrating that the worst-case approximation factor of WF is at least $K_{c}$. ∎
Theorem 1 establishes that $K_{c}$, the number of task groups in job $c$, is a lower bound for the approximation factor $\alpha_{\textsc{WF}}$ of WF. We now demonstrate that $K_{c}$ is also an upper bound for $\alpha_{\textsc{WF}}$, thereby tightly characterizing the approximation factor of WF.
Theorem 2.
For each newly arrived job $c$, the approximation factor of WF is at most $K_{c}$.
The proof is given in Appendix A.
Theorems 1 and 2 collectively establish that WF has a tight approximation ratio of $K_{c}$, where $K_{c}$ is the number of task groups in the incoming job $c$.
III-C Replica-Deletion
In this section, we introduce the Replica-Deletion (RD) heuristic, a task assignment strategy designed to minimize job completion time by iteratively balancing server workloads through the removal of redundant task replicas.
III-C1 Algorithm Design
We illustrate RD with an example, as shown in Fig. 5.
Initialization phase. When a new job $c$ arrives, each task $r\in\mathcal{T}_{c}$ is initially replicated $|\mathcal{S}^{r}$ times across all its available servers. For example, in Fig. 5(i), the blue task has three available servers (servers 1, 2, and 5), so it is replicated on these three servers.
Figure 5: How RD works. (i)–(ix) demonstrate how to delete task replicas. In this example, $\mu_{m}^{c}\equiv 1$ for all the servers $m\in\mathcal{M}$.
Deletion phase. After initialization, RD enters the deletion phase to iteratively remove redundant task replicas while ensuring the estimated busy times of participating servers remain as balanced as possible. The steps in each deletion iteration are as follows.
  1. 1.
    Identify target server(s). RD identifies the server(s) with the largest estimated busy time, referred to as the target server(s) ($m^{\star}$). For example, in Fig. 5(i), server 2 is the target server in the first iteration.
  2. 2.
    Delete task replicas. RD removes up to $\mu_{m^{\star}}^{c}$ task replicas from the target server to reduce its estimated busy time by one time slot. Among the tasks on the target server, the task with the largest number of replicas is prioritized for deletion. For example, in Fig. 5(i), server 2 hosts a red task with two replicas, a blue task with three replicas, and a pink task with two replicas. RD removes the blue task replica since it has the highest replication count.
  3. 3.
    Handle ties. If multiple tasks have the same maximum number of replicas, RD selects the server with the largest initial estimated busy time (i.e., before job $c$’s tasks are assigned) as the target. For example, in Fig. 6, servers 1 and 5 both have three replicas of the blue and green tasks, respectively. RD deletes the blue task replica from server 1 because server 1 has a higher initial estimated busy time.
  4. 4.
    Repeat until termination. The deletion phase continues until all tasks in the target servers have only one replica. At this point, the completion time of job $c$ cannot be further reduced. For example, in Fig. 5(viii), RD terminates the deletion phase when all tasks in servers 1 and 5 have no other replicas.
Finalization phase. In the finalization phase, RD removes any remaining redundant task replicas to ensure that each task is assigned to only one server. This is done by continuing the replica removal process, focusing on servers with the highest estimated busy times. For example, in Fig. 5(ix), RD deletes the yellow and green task replicas from server 4 to balance workloads among the remaining participating servers.
Figure 6: A target server with a larger initial estimated busy time has a task replica removed.
RD differs from WF in its global perspective on task assignment. RD evaluates all available servers when deciding which task replica to delete, enabling it to achieve a more balanced workload distribution across the entire system. In contrast, WF operates locally by dividing tasks into groups and balancing workloads within each group during each iteration. While RD’s global approach often yields better performance, it also increases the computational complexity of the assignment process.
III-C2 Complexity Analysis
In the implementation of RD, we use a priority queue to manage all servers, prioritizing them by their estimated busy times and breaking ties using their initial estimated busy times. Additionally, each server $m$ maintains its own priority queue to manage the tasks assigned to it, prioritizing tasks by their number of replicas. In the worst-case scenario, each task is replicated across all servers, requiring $\mathcal{O}(|\mathcal{T}_{c}|\cdot M)$ deletions. Each deletion operation affects the priority queues of all servers containing replicas of the deleted task. Updating the priority queue of up to $M$ servers requires $\mathcal{O}(M\cdot\log|\mathcal{T}_{c}|)$ time per deletion. Thus, the total time complexity of RD is:
$\mathcal{O}\Big{(}M^{2}\cdot|\mathcal{T}_{c}|\cdot\log|\mathcal{T}_{c}|\Big{)}.$
We experimentally compare the performance of RD with WF and other algorithms in Sec. V. Theoretical performance analysis of RD is left as future work.
IV Algorithms with Job Reordering
This section considers an extended scenario in which outstanding job tasks queued on servers can be prioritized and reordered. By adjusting the execution order of outstanding jobs and reassigning their remaining tasks, the average completion time of jobs can be further reduced.
Inspired by [6] and [9], we adopt a similar approach to derive the job execution order, which emulates a shortest-remaining-time-first (SRTF) policy for distributed job executions. When a new job $c$ arrives, reordering is triggered. Let $\mathcal{O}_{c}$ denote the set of outstanding jobs after the arrival of job $c$, and let $\mathcal{Q}_{c}$ represent the reordered list of outstanding jobs. The goal is to iteratively retrieve jobs from $\mathcal{O}_{c}$ and insert them into $\mathcal{Q}_{c}$ in a specific sequence until $|\mathcal{Q}_{c}|=|\mathcal{O}_{c}|$.
[…]
Input: Online arriving jobs $1,2,...,c,...$, server capacities $\{\mu_{m}^{c}\}_{m,c}$
Output: Assignment solution for each task of each arriving job
1 while a new job $c$ arrives do
2 $\mathcal{O}_{t}\leftarrow\mathcal{O}_{t}\cup{c}$
3 $\mathcal{Q}_{t}\leftarrow\emptyset$
4   Initialize estimated busy times of servers: $\forall m\in\mathcal{M}$, $b_{m}^{c}\leftarrow 0$
5 while $|\mathcal{Q}_{c}|<|\mathcal{O}_{c}|$ do
6    $l\leftarrow\texttt{nullptr}$
7    for each job $c’\in\mathcal{O}_{t}\backslash\mathcal{Q}_{t}$ do
8         Get unprocessed tasks of job $c’$ ’ and calculate $\Phi_{c’}^{-}$ using (8) and (9)
9       if $l==\texttt{nullptr}$ or $\Phi_{c’}^{-}<\Phi_{l}$ then
10            Solve the task assignment problem $\mathcal{P}$ for job $c’$ ’ using WF to calculate $\Phi_{c’}$
11          if $l==\texttt{nullptr}$ or $\Phi_{c’}<\Phi_{l}$ then
12               Derive the task assignment for $c’$ ’ using WF
13             $l\leftarrow c’$ ’
14
15       else
          break // early-exit
16
17
18    $\mathcal{Q}_{c}\leftarrow\mathcal{Q}_{c}\cup\{l\}$
19      For each $m\in\mathcal{M}$, update the estimated busy time of server $m$
20
Algorithm 3 OCWF-ACC
We name the algorithm as OCWF-ACC (accelerated order-conscious WF). Algorithm 3 presents the details of OCWF-ACC. Note that WF can be replaced by other task assignment algorithms.
V Experimental Results
We conduct extensive simulations to verify the performance and efficiency of the proposed algorithms. We first present the simulation settings. Then, we show and discuss the results.
V-A Experimental Setup
Figure 7: Average job completion time, computation overheads, and CDF of job completion times under 25% system utilization for different algorithms.
Figure 8: Average job completion time, computation overheads, and CDF of job completion times under 50% system utilization for different algorithms.
Figure 9: Average job completion time, computation overheads, and CDF of job completion times under 75% system utilization for different algorithms.
Job Traces. We used the Alibaba cluster trace dataset [10] for simulation. Specifically, a segment from the dataset batch_task.csv in cluster-trace-v2017 was selected, containing 250 jobs with a total of 113,653 tasks. Job arrivals and task durations were derived from the timestamps of recorded task events. The inter-arrival times of jobs were scaled to simulate system utilizations of 25%, 50%, and 75%. The default number of servers was 100, with each server’s processing capacity for a given job randomly generated between 3 and 5 by default.
Available Servers. Each job entry in the trace corresponds to a task group. On average, each job contains 5.52 task groups. Data inputs for task groups were distributed across servers according to a Zipf distribution. For each task group, a random permutation of all servers was generated, and the task group was associated with the $i$-th server in the permutation with a probability proportional to $\frac{1}{i^{\alpha}}$. Here, $\alpha\in[0,2]$ is the Zipf skew parameter, where $\alpha=0$ indicates a uniform distribution. For a task group associated with server $m$, servers $m,m+1,\dots,m+p-1$ were chosen as its available servers, where $p$ was randomly generated between 8 and 12 by default.
Algorithms. We implemented six algorithms using DOcplex: [^2] https://ibmdecisionoptimization.github.io/docplex-doc/index.html NLIP, OBTA, WF, RD, OCWF, and OCWF-ACC.
  • NLIP solves the non-linear program $\mathcal{P}$ directly for each job, without narrowing the search space or dividing it into subranges, as in OBTA.
  • OBTA refines NLIP by reducing the search space for $\Phi_{c}$, improving efficiency while maintaining optimality (see Sec. III-A).
  • WF assigns tasks approximately in a water-filling manner to achieve a good balance between performance and overhead (see Sec. III-B).
  • RD employs a task assignment strategy based on an initialization-then-deletion process, aiming to iteratively balance server workloads (see Sec. III-C).
  • OCWF reorders jobs to prioritize those with shorter remaining times but does not employ the early-exit optimization used in OCWF-ACC.
  • OCWF-ACC extends OCWF with the early-exit technique to reduce computational overhead (see Sec. IV).
Metrics. We use the average job completion time of all jobs to measure performance and the computation overhead of each algorithm to measure efficiency.
The code is available online. [^3] https://github.com/hliangzhao/taos
V-B Performance Evaluation
We compare the performance and efficiency of the algorithms by varying the Zipf parameter $\alpha$ from 0 to 2 and the system utilization from 25% to 75%. Figs. 7-9 present the results. In these figures, the right y-axis in the first subfigure indicates the average job completion time, while the left y-axis shows the log-scaled average computation overhead per job arrival. The other four subfigures visualize the cumulative distribution of job completion times for different $\alpha$ values. The key findings are summarized as follows.
Compared with NLIP, OBTA achieves a significant improvement in efficiency. From Figs. 7-9, we observe that OBTA and NLIP have fairly close performance of job completion time since both are theoretically optimal for balanced task assignment. By narrowing the search space and dividing it into subranges, OBTA reduces the computation overhead by nearly half compared to NLIP. This verifies our contribution in Sec. III-B. The notable reduction in computational overhead without sacrificing performance demonstrates the scalability of OBTA. This is especially relevant for large-scale systems where tasks arrive dynamically, and the computational cost of solving NLIP becomes a bottleneck. OBTA, by efficiently pruning the search space, provides a more practical approach for real-world applications, offering significant gains in execution speed while maintaining optimal task assignment.
WF closely approximates balanced task assignment with extremely low overhead. WF performs close to OBTA and NLIP at almost all percentiles of job completion times, and its computation overhead is two orders of magnitude lower than OBTA’s. The ultra-low overhead offers a significant advantage and facilitates job reordering. The lightweight nature of WF makes it particularly appealing in scenarios where the frequency of task arrivals is high, and real-time decisions are necessary. Although WF sacrifices some degree of optimality, its approximation is sufficiently close to the theoretical optima for most practical purposes. The low computational cost not only accelerates the scheduling process but also allows it to be integrated effectively into more complex frameworks like OCWF and OCWF-ACC.
RD generally produces better task assignments than WF. Our results show that RD generally outperforms WF. This observation is detailed in Table III, which presents the performance of each algorithm when $\alpha=2$ and system utilization is at 75%, as shown in the first figure of Fig. 10. However, it is noteworthy that RD has a higher computation overhead than WF, which aligns with our complexity analysis. The time complexity of RD is quantified as $\mathcal{O}(M^{2}\cdot n\log n)$, which is higher than WF’s $\mathcal{O}(KM\cdot\log n)$, where $M$ is the number of servers, $n$ is the number of tasks, and $K$ is the number of task groups in a job. RD’s global perspective on load balancing allows it to achieve better overall performance than WF in terms of job completion times. However, its relatively higher computational overhead makes it less suitable for latency-sensitive applications. Despite this, RD’s balancing strategy is particularly effective in systems with heterogeneous resources or non-uniform task distributions, where WF might underperform due to its localized approach.
TABLE III: Average job completion time vs. #available servers.
Algorithms Number of available servers
4 6 8 10 12 Aver.
OBTA 12551 5814 4217 3609 3160 5870
NLIP 12468 5924 4394 3547 3164 5899
WF 12872 6125 4408 3617 3188 6042
RD 12857 5987 4261 3595 3150 5970
OCWF 1699 958 765 712 657 958
OCWF-ACC 1699 958 765 712 657 958
OCWF-ACC and OCWF are robust to skewness in data availability. As shown in Figs. 7-9, when the skew parameter $\alpha$ changes from 0 to 2, the average job completion time achieved by FIFO-based algorithms (NLIP, OBTA, WF, and RD) shows a clear upward trend. In contrast, OCWF and OCWF-ACC do not exhibit a significant increase in average job completion time. This demonstrates that reordering and reallocation are crucial for handling skewness in data availability. The robustness of OCWF and OCWF-ACC against skewed data distributions underscores the importance of reordering in mitigating the impact of resource contention. By dynamically adjusting job priorities based on estimated completion times, these algorithms effectively balance workloads across servers, even under highly non-uniform conditions.
OCWF-ACC significantly accelerates OCWF. The computation overhead of OCWF-ACC is only about half of OCWF’s. This verifies the advantage of the early-exit technique. The early-exit optimization not only reduces the time required for job reordering but also enhances the scalability of OCWF-ACC. This makes it a practical choice for systems with stringent latency requirements, as it balances the trade-off between computational efficiency and scheduling effectiveness.
Figure 10: Average job completion time, and CDF of job completion times when $\alpha=2$ under 75% system utilization for different numbers of available servers.
Figure 11: Average job completion time, and CDF of job completion times when $\alpha=2$ under 75% system utilization for different processing capacities.
Figs. 7-9 illustrate that job completion times increase with system utilization. The performance trends discussed above are consistently observed across different levels of system utilization. Higher utilization levels amplify resource contention, which tests the scalability and robustness of the scheduling algorithms. While FIFO-based approaches struggle under these conditions, the reordering and reallocation capabilities of OCWF and OCWF-ACC enable them to maintain stable performance, even at high utilization.
Figs. 10-11 show the job completion times for varying numbers of available servers and profiled processing capacities, respectively. These experiments are conducted with $\alpha=2$ and a system utilization of 75%, as we are particularly interested in the performance when resource contention is high. From Fig. 10, we observe that a larger number of available servers decreases job completion times. More available servers provide greater flexibility in task assignment, allowing tasks to be distributed in a manner that better balances the load among servers and reduces job completion times. Fig. 11 demonstrates that increasing processing capacity also reduces job completion times. This is expected, as higher processing capacities enable more tasks to run in parallel, thereby decreasing job completion times. Notably, the relative performance of the algorithms remains largely unchanged across different numbers of available servers and processing capacities.
V-C Addition Evaluation on Algorithm Efficiency
In this section, we evaluate the effectiveness of OBTA’s two key optimization components: solution-space narrowing and piecewise linearization, by comparing OBTA with its variants. To isolate and evaluate the impact of each optimization strategy in OBTA, we implemented and compared it with three alternative approaches:
  • LIP: A direct solution to the linearized version of the original problem without any search-space reduction. The objective of the linearized problem $\mathcal{P}^{\prime}$ is the same as $\mathcal{P}$, but the first set of non-linear constraints in (6) are transformed into the following ones:
    $\forall m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}:\left\{\begin{array}[]{l}\sum_{k\in\mathcal{K}_{c}}n_{m}^{k}\leq z_{m},\\ z_{m}\geq\Phi_{c}-b_{m}^{c},\\ z_{m}\geq 0,\\ z_{m}\leq\Phi_{c}-b_{m}^{c}+\lambda\cdot y_{m,1},\\ z_{m}\leq\lambda\cdot y_{m,2},\\ y_{m,1}+y_{m,2}\leq 1,\\ y_{m,1},y_{m,2}\in\{0,1\},\end{array}\right.$ (27)
    where $\lambda$ is a sufficiently large number.
  • OBTA-N: A variant of OBTA that removes the piecewise linearization component and retains only the solution-space narrowing mechanism.
  • OBTA-P: A variant of OBTA that removes the solution-space narrowing step and applies full-range piecewise linearization.
TABLE IV: Computation overheads of algorithms to complete the scheduling of all jobs (in second).
Algorithms Zipf Parameter $\alpha$
$\alpha=0$ $\alpha=0.5$ $\alpha=1$ $\alpha=2$
OBTA 14.58 15.75 22.31 40.54
NLIP 21.60 24.42 33.89 48.17
LIP 17.79 17.89 27.82 42.06
OBTA-N 18.59 18.97 25.49 43.96
OBTA-P 131.26 133.94 136.59 85.42
WF 10.12 9.63 16.16 37.25
RD 12.48 10.98 15.93 37.79
TABLE V: Average job completion time achieved by algorithms.
Algorithms Zipf Parameter $\alpha$
$\alpha=0$ $\alpha=0.5$ $\alpha=1$ $\alpha=2$
OBTA 694 823 1517 7081
NLIP 709 813 1494 7164
LIP 682 798 1571 7183
OBTA-N 704 831 1518 7151
OBTA-P 703 818 1548 7213
WF 755 861 1569 7270
RD 772 852 1567 7252
We evaluated all four algorithms: OBTA, LIP, OBTA-N, and OBTA-P, alongside WF, RD, and NLIP under the experimental settings demonstrated in Sec. V-A (the system utilization is set to 75% by default). As shown in Tables IV and V, in terms of solution quality, all algorithms perform similarly, indicating that the linearization and search-space reduction do not compromise optimality. However, in terms of computational efficiency: WF $>$ RD $>$ OBTA $>$ LIP $>$ OBTA-N $>$ NLIP $>$ OBTA-P, where ‘$>$’ indicates lower computational cost (i.e., faster execution). These results demonstrate that both components of OBTA contribute to its efficiency. In particular, solution-space narrowing (OBTA-N) achieves the lowest runtime among all variants, suggesting that this component has the most significant impact on reducing computational overhead. Meanwhile, the combination of both strategies in OBTA achieves the best balance between efficiency and stability across different workloads.
V-D Robustness Under Large-Scale and Heterogeneous Settings
To further assess scalability and robustness, we tested the algorithms under larger-scale scenarios and heterogeneous server environments.
First, we increased the input size to include 500 jobs, 256 servers, and 173,848 tasks, significantly expanding the solution space. The system utilization is set to 75% by default. As shown in Table VI, the relative performance trends remain consistent, confirming that OBTA and its variants scale effectively with increasing problem size.
TABLE VI: Large-scale experiments with 75% system utilization and Zipf parameter $\alpha$ being 2.
Algorithms OBTA NLIP LIP OBTA-N OBTA-P WF RD OCWF OCWF-ACC
Overhead (s) 89 96 104 101 221 78 77 6538 3118
Average job completion time 8639 8764 9059 8716 8878 9281 8873 1189 1234
Next, we evaluated the algorithms under various server capacity distributions to simulate heterogeneous computing environments. Specifically, we considered uniform, power-law, log-normal, and weighted distributions. These models reflect real-world diversity in server capabilities. Tables VII and VIII summarize the average job completion time and computational overheads under these distributions. Across all settings, the relative rankings of algorithmic efficiency remain stable, demonstrating that the performance improvements achieved by OBTA are not sensitive to infrastructure heterogeneity. These experiments confirm that OBTA maintains its efficiency advantage even when scaling to large workloads, and the proposed algorithms perform consistently well under diverse server capacity distributions, showing their robustness to system heterogeneity.
TABLE VII: Average job completion time achieved by algorithms under different server capacity distributions.
Algorithms Types of Server Capacity Distribution
Uniform Power-law Log-normal Weighted
OBTA 6713 6530 7667 8872
NLIP 6935 6803 7833 8987
LIP 6865 6433 7878 9062
OBTA-N 7076 6778 7602 8899
OBTA-P 6821 6549 7622 8812
WF 6966 6599 7673 8846
RD 7071 6819 7691 8856
OCWF 1172 1118 1634 1674
OCWF-ACC 1222 1176 1634 1674
TABLE VIII: Computation overheads of algorithms under different server capacity distributions (in second).
Algorithms Types of Server Capacity Distribution
Uniform Power-law Log-normal Weighted
OBTA 36.61 34.27 50.11 54.37
NLIP 39.97 39.35 57.73 57.08
LIP 40.75 36.88 52.26 56.03
OBTA-N 42.09 38.06 51.96 56.84
OBTA-P 85.23 75.35 99.38 103.93
WF 32.63 29.22 43.16 45.95
RD 34.76 31.67 46.03 47.05
OCWF 578.82 725.53 769.46 780.93
OCWF-ACC 410.00 399.36 512.85 508.66
To test the sensitivity of our approaches to how tasks map to available servers, we modified the distribution of server availability. Specifically, we replaced the original Zipf-based assignment with log-normal, exponential, and uniform distributions. As shown in Tables IX and X, the experimental results further support our initial conclusions, indicating that the performance trends of our algorithms are largely insensitive to the specific distribution patterns.
TABLE IX: Average job completion time achieved by algorithms under different available server distributions.
Algorithms Types of Available Server Distribution
Zipf Log-normal Exponential Uniform
OBTA 6353 800 783 767
NLIP 6513 803 780 745
LIP 6300 790 804 784
OBTA-N 6410 770 777 747
OBTA-P 6486 816 779 740
WF 6529 825 804 771
RD 6750 818 806 817
OCWF 1197 345 345 311
OCWF-ACC 1204 339 342 333
TABLE X: Computation overheads of algorithms under different available server distributions (in second).
Algorithms Types of Available Server Distribution
Zipf Log-normal Exponential Uniform
OBTA 32.30 17.07 14.82 15.80
NLIP 35.93 22.85 21.25 19.60
LIP 36.14 19.94 17.29 18.37
OBTA-N 36.17 20.17 18.61 19.48
OBTA-P 76.16 117.04 110.96 107.29
WF 27.64 8.75 8.39 10.35
RD 31.21 10.21 9.89 9.98
OCWF 731.56 569.83 574.69 559.44
OCWF-ACC 408.20 419.25 432.90 426.22
VI Related Work
Job scheduling has been extensively studied from both theoretical [11, 12, 13, 14, 15, 16, 17, 18, 19] and system-level perspectives [20, 21, 22, 23, 24, 25]. Theoretical works often formulate job completion time minimization as combinatorial, constrained optimization problems, addressing them using a variety of techniques, particularly approximation algorithms. These studies provide foundational insights into optimality and computational tractability, often emphasizing complexity bounds and performance guarantees.
A critical system-level consideration in distributed scheduling is data locality, which emphasizes assigning tasks close to where the required data chunks reside to minimize communication overhead. Rooted in the MapReduce paradigm [5], this principle has driven numerous scheduling algorithms for both homogeneous and heterogeneous systems [26, 27, 28, 8]. These approaches typically aim to balance data locality against other performance goals such as load balancing [29], throughput [30], delay [28, 31], fairness [7], and job completion time [32, 6, 33, 34, 35, 36].
Within this context, scheduling strategies that explicitly preserve data locality in multi-task, online job scenarios have received increased attention [6, 37, 7, 8, 9]. For instance, Hung et al. [6] proposed SWAG, a workload-aware greedy algorithm that schedules jobs by estimating the completion times of all outstanding jobs. SWAG was the first to formalize job reordering in distributed environments with data locality constraints.
TABLE XI: Comparison of related works for distributed job executions.
Related Work Online Data Locality Multi-task Job Theoretical Analysis Real Trace Evaluation
SWAG [6]
BTAaJ / ATA-Greedy [9]
SG-PBFS / EG-SJF [34, 36]
MTD-DHJS [35]
This Work
Building on this, Guan et al. [9] proposed the BTAaJ algorithm, which significantly improved upon SWAG by utilizing a well-constructed flow network to assign tasks for newly arrived jobs. BTAaJ groups the tasks of a job based on their available servers, achieving load balancing by minimizing a control parameter $C$. Additionally, Guan et al. proposed the ATA-Greedy algorithm, which enhances scheduling performance by reordering outstanding jobs. This reordering process builds on the approach in [6], incorporating the control parameter $C$ to refine completion time estimations. Our work extends these efforts by introducing algorithms that offer both improved performance and reduced computational overhead, making significant advancements over [6] and [9].
Other dimensions of distributed job scheduling, such as fairness and data movement, have also been explored. For instance, Guan et al. [7] extended the concept of max-min fairness from single-machine settings to distributed job executions over multiple servers. Their Aggregate Max-min Fairness policy is designed to be Pareto-efficient, envy-free, and strategy-proof, providing equitable resource allocation in distributed systems. Beaumont et al. [8] addressed the Comm-Oriented problem, which optimizes the makespan while minimizing data movement. However, their focus on single-job scenarios limits its applicability to multi-job online scheduling, which is the primary concern of our work. Several recent works have also proposed heuristic scheduling algorithms for cloud environments. For instance, Banerjee et al. [35] proposed MTD-DHJS, which combines Johnson sequencing and Round Robin to reduce makespan on a fixed three-server architecture. However, their model assumes batch-style job processing and lacks support for task-level data locality, complex job structures, or online arrivals. Similarly, Murad et al. [34, 36] developed SG-PBFS and EG-SJF, which enhance traditional priority-rule (PR) scheduling with gap-based backfilling to improve flow time and delay. These methods rely on metadata such as CPU time and arrival time for decision-making, but diverge from our work in both modeling assumptions and distributed settings. In particular, they do not handle replicated input data, multi-task jobs, or theoretical performance guarantees.
VII Conclusion
In this paper, we explored the task assignment problem in the context of online distributed job executions, with data locality preserved. Our contributions began with formulating the task assignment problem as a non-linear integer program. We developed the OBTA algorithm, known for its efficiency in reducing computational overhead by narrowing the search space of potential solutions. Additionally, we extended the WF heuristic, providing an in-depth, nontrivial tight analysis of its approximation factor, and introduced the innovative RD algorithm, which demonstrates enhanced performance. Our investigation into job reordering, particularly with the implementation of the novel early-exit technique, effectively optimized average job completion times while maintaining low computational overheads. The effectiveness of all proposed algorithms was validated through extensive trace-driven simulations. Both OBTA and the early-exit technique showed substantial reductions in computational overhead. Meanwhile, WF emerged as a practical, lower-overhead alternative for large-scale scenarios. Despite its greater computational intensity compared to WF, RD exhibited superior performance of job completion time. Looking ahead, integrating these algorithms into diverse computational settings and adapting them to the evolving dynamics of distributed job executions hold potential for further enhancing their applicability.
Appendix A Proof of Theorem 2
We need to prove that $\textsc{WF}(I)\leq K_{c}\cdot\textsc{OPT}(I)$ for any possible arrival instance $I$. For simplicity, we omit $I$ in the notation and use ALG to represent both the algorithm and the estimated completion time of job $c$ under that algorithm. ALG can refer to OPT, WF, $\textsc{WF}_{k}$, or $\textsc{OPT}_{k}$ (the last two notations will be introduced shortly).
[…]
$\textsc{WF}_{k}=\max_{m\in\bigcup_{k^{\prime}\leq k}\Omega_{c}^{k^{\prime}}}b_{m}^{c}(k),$ (28)
and for each $k<K_{c}$,
$\textsc{WF}_{k+1}=\max\left\{\textsc{WF}_{k},\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\right\}.$ (29)
Recall that $b_{m}^{c}(k)$ is updated by (13). Fig. 12 visualizes $\textsc{WF}_{k}$.
Figure 12: Visualization of $\textsc{WF}_{k}$. Note that $\textsc{WF}_{k}=\textsc{WF}_{k-1}$ for the constructed example.
Since $\mathcal{T}_{c}^{1}\subseteq\mathcal{T}_{c}$, $\textsc{WF}_{1}\leq\textsc{OPT}$. Note that $\textsc{WF}=\textsc{WF}_{K_{c}}$. If we can prove that for each $k<K_{c}$,
$\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT},$ (30)
then we have
$\textsc{WF}=\textsc{WF}_{1}+\sum_{k<K_{c}}(\textsc{WF}_{k+1}-\textsc{WF}_{k})\leq K_{c}\cdot\textsc{OPT},$ (31)
which is exactly our target. In the following, we will show that (30) indeed holds. To do so, we exploit the optimal assignment of the tasks of a job containing the tasks $\mathcal{T}_{c}^{k+1}$ only. Similar to $\textsc{WF}_{k+1}$, we denote by $\textsc{OPT}_{k+1}$ the number of time slots required to complete the processing of tasks in $\mathcal{T}_{c}^{k+1}$. It is easy to see that
$_{k+1}\leq\textsc{OPT}$ OPT (32)
holds for each $k\in\mathcal{K}_{c}$, since $\mathcal{T}_{c}^{k+1}\subseteq\mathcal{T}_{c}$.
To prove (30), we consider different cases of the assignment of tasks in $\mathcal{T}_{c}^{k+1}$ by WF separately.
Case I. $\mathcal{S}_{c}^{k+1}\cap\bigcup_{k^{\prime}\leq k}\Omega_{c}^{k^{\prime}}=\emptyset$. In this case, the available servers for group $k+1$ do not overlap with the participating servers for groups $1,\ldots,k$. Thus, by the definition of WF, such an assignment must be optimal if a job contains only the tasks $\mathcal{T}_{c}^{k+1}$. Hence, by (32), we have
$\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)=\textsc{OPT}_{k+1}\leq\textsc{OPT}.$ (33)
  • As visualized in Fig. 13(i), if $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$, by (29), we have $\textsc{WF}_{k+1}=\textsc{WF}_{k}$. Thus,
    $\textsc{WF}_{k+1}-\textsc{WF}_{k}=0<\textsc{OPT}.$ (34)
  • As visualized in Fig. 13(ii), if $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$, by (29) and (33), we have
    $\textsc{WF}_{k+1}-\textsc{WF}_{k}$ $<\textsc{WF}_{k+1}$
    $=\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)$
[…]
Figure 13: Visualization of Case I. In (i), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$. In (ii), $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$.
Case II. $\mathcal{S}_{c}^{k+1}\cap\bigcup_{k^{\prime}\leq k}\Omega_{c}^{k^{\prime}}\neq\emptyset$, i.e., the available servers of group $k+1$ overlap with the participating servers of lower-indexed groups.
If $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)\leq\textsc{WF}_{k}$, we again have $\textsc{WF}_{k+1}=\textsc{WF}_{k}$ by (29) and thus, $\textsc{WF}_{k+1}-\textsc{WF}_{k}=0<\textsc{OPT}$.
If $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)>\textsc{WF}_{k}$, we have $\textsc{WF}_{k+1}>\textsc{WF}_{k}$. In this case, by definition, each participating server of groups $1,\ldots,k$ must also be a participating server of group $k+1$ by WF (because the estimated busy times of all these servers are smaller than $\max_{m\in\Omega_{c}^{k+1}}b_{m}^{c}(k+1)$).
For each $m\in\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1}$ (i.e., server $m$ is available to group $k+1$ but does not participate in $\mathcal{T}_{c}^{k+1}$’s assignment), it can be shown that:
  1. 1.
    $b_{m}^{c}(k)\geq\max_{m^{\prime}\in\Omega_{c}^{k+1}}b_{m^{\prime}}^{c}(k+1)$, and
  2. 2.
    by WF, server $m$ does not participate in the assignment of groups $1,\ldots,k$.
Property 1) follows directly from the definition of WF (otherwise, server $m$ must be a participating server of group $k+1$). Property 2) follows from the above observation that each participating server of groups $1,\ldots,k$ must also be a participating server of group $k+1$ by WF. By properties 1) and 2), for each server $m\in\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1}$, the initial estimated busy time $b_{m}^{c}(0)$ satisfies
$b_{m}^{c}(0)=b_{m}^{c}(k)\geq\max_{m^{\prime}\in\Omega_{c}^{k+1}}b_{m^{\prime}}^{c}(k+1).$ (36)
With this result, we prove $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$ by comparing $\Omega_{c}^{k+1}$ and $\beth_{c}^{k+1}$, where $\beth_{c}^{k+1}\subseteq\mathcal{S}_{c}^{k+1}$ is the set of participating servers for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$.
Case II-A. $\beth_{c}^{k+1}\cap(\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1})\neq\emptyset$, i.e., at least one server in $\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1}$ is chosen for assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$. Taking any server $m\in\beth_{c}^{k+1}\cap(\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1})$, we have
[…]
As a result, $\textsc{WF}_{k+1}-\textsc{WF}_{k}<\textsc{WF}_{k+1}<\textsc{OPT}$.
Case II-B. $\beth_{c}^{k+1}\cap(\mathcal{S}_{c}^{k+1}\backslash\Omega_{c}^{k+1})=\emptyset$. It follows that $\beth_{c}^{k+1}\subseteq\Omega_{c}^{k+1}$, which implies
$\beth_{c}^{k+1}=\Omega_{c}^{k+1}\cap\beth_{c}^{k+1}.$ (38)
We divide the servers in $\Omega_{c}^{k+1}$ into two disjoint sets:
$\Omega_{c,\leq}^{k+1}:=\left\{m\in\Omega_{c}^{k+1}\mid b_{m}^{c}(k)\leq\textsc{WF}_{k}\right\},$ (39)
$\Omega_{c,>}^{k+1}:=\left\{m\in\Omega_{c}^{k+1}\mid b_{m}^{c}(k)>\textsc{WF}_{k}\right\}.$ (40)
For each $m\in\Omega_{c,>}^{k+1}$, if server $m$ is a participating server in the assignment of any group $1,\ldots,k$ by WF, we have
[…]
giving a contradiction. Thus, server $m$ is not a participating server in the assignment of groups $1,\ldots,k$ by WF. As a result, $b_{m}^{c}(k)=b_{m}^{c}(0)$ holds for each $m\in\Omega_{c,>}^{k+1}$.
Case II-B1. $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}=\emptyset$. In this case, we have $\beth_{c}^{k+1}\subseteq\Omega_{c,\leq}^{k+1}$. In $\textsc{OPT}_{k+1}$, $\textsc{OPT}_{k+1}$ time slots of all the participating servers $\beth_{c}^{k+1}$ can accommodate all the tasks of $\mathcal{T}_{c}^{k+1}$. Thus,
$|\mathcal{T}_{c}^{k+1}|\leq\textsc{OPT}_{k+1}\cdot\sum_{m\in\beth_{c}^{k+1}}\mu_{m}^{c}.$ (42)
On the other hand, by the definition of WF, $(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)$ time slots of all the participating servers $\Omega_{c,\leq}^{k+1}$ are not adequate for all the tasks of $\mathcal{T}_{c}^{k+1}$. Hence,
$\left(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1\right)\cdot\sum_{m\in\Omega_{c,\leq}^{k+1}}\mu_{m}^{c}<|\mathcal{T}_{c}^{k+1}|.$ (43)
Combining (42) and (43) and noting that $\beth_{c}^{k+1}\subseteq\Omega_{c,\leq}^{k+1}$, we obtain $\textsc{WF}_{k+1}-\textsc{WF}_{k}-1<\textsc{OPT}_{k+1}$. By (32), we have $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$.
Case II-B2. $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}\neq\emptyset$. In this case, at least one server in $\Omega_{c,>}^{k+1}$ is a participating server for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$. Hence, all the servers in $\Omega_{c,\leq}^{k+1}$ must also be participating servers for the assignment of $\mathcal{T}_{c}^{k+1}$ by $\textsc{OPT}_{k+1}$. In other words, $\Omega_{c,\leq}^{k+1}\subseteq\beth_{c}^{k+1}$.
Figure 14: Visualization of Case II-B2.
In $\textsc{OPT}_{k+1}$, $\textsc{OPT}_{k+1}$ time slots of all the participating servers $\beth_{c}^{k+1}$ can accommodate all the tasks of $\mathcal{T}_{c}^{k+1}$ and the outstanding tasks of previous jobs queued at the servers in $\beth_{c}^{k+1}\cap\Omega_{c,>}^{k+1}$. Thus,
$|\mathcal{T}_{c}^{k+1}|+\sum_{m\in\Omega_{c,>}^{k+1}\cap\beth_{c}^{k+1}}b_{m}^{c}(0)\cdot\mu_{m}^{c}\leq\textsc{OPT}_{k+1}\cdot\sum_{m\in\beth_{c}^{k+1}}\mu_{m}^{c}.$ (44)
On the other hand, by the definition of WF, $(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)$ time slots of all the participating servers in $\Omega_{c}^{k+1}\cap\beth_{c}^{k+1}$ are not adequate for all the tasks of $\mathcal{T}_{c}^{k+1}$ and the backlogs beyond $\textsc{WF}_{k}$ (see the black box in Fig. 14 for an illustration). Hence,
$(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)\cdot\sum_{m\in\beth_{c}^{k+1}}\mu_{m}^{c}$
=(38)${=}(\textsc{WF}_{k+1}-\textsc{WF}_{k}-1)\cdot\sum_{m\in\Omega_{c}^{k+1}\cap\beth_{c}^{k+1}}\mu_{m}^{c}$
$\qquad<|\mathcal{T}_{c}^{k+1}|+\sum_{m\in\Omega_{c,>}^{k+1}\cap\beth_{c}^{k+1}}(b_{m}^{c}(0)-\textsc{WF}_{k})\cdot\mu_{m}^{c}.$ (45)
Note that the right side of (45) is no greater than the left side of (44). Combining (44) with (45), we obtain $\textsc{WF}_{k+1}-\textsc{WF}_{k}-1<\textsc{OPT}_{k+1}$. By (32), we have $\textsc{WF}_{k+1}-\textsc{WF}_{k}\leq\textsc{OPT}$.
With the above analysis for Case I and Case II, we have shown that (30) indeed holds.